### v2

# Theoretical Foundations of Prompt Engineering: From Heuristics to Expressivity

Abstract
Prompts can switch a model’s behavior even when the weights are fixed, yet this phenomenon is rarely treated as a clean theoretical object rather than a heuristic. We study the family of functions obtainable by holding a Transformer backbone fixed as an executor and varying only the prompt. Our core idea is to view the prompt as an externally injected program and to construct a simplified Transformer that interprets it to implement different computations. The construction exposes a mechanism-level decomposition: attention performs selective routing from prompt memory, the FFN performs local arithmetic conditioned on retrieved fragments, and depth-wise stacking composes these local updates into a multi-step computation. Under this viewpoint, we prove a constructive existential result showing that a single fixed backbone can approximate a broad class of target behaviors via prompts alone. The framework provides a unified starting point for formalizing trade-offs under prompt length/precision constraints and for studying structural limits of prompt-based switching, while remaining distinct from empirical claims about pretrained LLMs.
Machine Learning, ICML
1 Introduction
Prompts often function like a switch: with the same underlying model, changing only the prompt can effectively change the task. In practice, modifying the prompt alone can drastically alter the output format, the style of reasoning, and even characteristic failure modes. Yet prompts are still largely discussed at the level of ”tips and tricks,” and it remains difficult to state mathematically why such switching is possible, how far it can go, and when or why it breaks. This gap has made it hard to systematize prompt engineering as a theoretical subject, despite its practical impact.
This paper reframes the question from ”what can a model do?” to ”what can a prompt change in a fixed model?” The core idea is to view a prompt not merely as an instruction string but as an external parameter (a program input) interpreted by a fixed Transformer backbone. This viewpoint lets us treat, as a single object, the range of behaviors achievable by varying prompts, and to restate ”why/how far/when does it fail?” in the language of expressivity and structural constraints.
Building on this perspective, we prove a foundational possibility result: there exists a single fixed Transformer that, with suitably chosen prompts, can approximate any continuous target behavior on a compact input domain to arbitrarily small error. Importantly, we do not leave this as a purely declarative statement. We give a constructive account of the internal roles that enable switching and make them reusable as lemmas: attention provides selective routing from prompt memory, the FFN provides local arithmetic using retrieved fragments, and depth-wise repetition provides composition that accumulates multi-step computation across blocks. Our contribution is not an empirical claim about pretrained LLMs, nor an algorithm for finding effective prompts under realistic constraints; rather, it provides a precise framework that makes such follow-up questions well-posed, including trade-offs under length and precision constraints, constraint-driven limits, and learnability or search procedures for finding good prompts.
Contributions.
  1. 1.
    Prompts as a theoretical object. We formalize the prompt-induced behaviors of a fixed model as a single object, enabling an expressivity-based view of prompt design.
  2. 2.
    An existence result for fixed-backbone prompt switching. We show that a single fixed Transformer can approximate a broad target class by changing only the prompt, without task-specific weight updates.
  3. 3.
    A structural explanation of why switching works. We explain switching via a lemma-based decomposition into routing (selective readout), arithmetic (local computation), and composition (multi-step accumulation) across blocks.
  4. 4.
    A roadmap for quantitative and algorithmic questions. We identify principled directions including length/precision/norm trade-offs, limits for finite hard prompts, constraints induced by pretraining and architecture, and the learnability of effective prompts under practical optimization.
2 Related Work
2.1 Universal Approximation and Expressivity Foundations
Universal approximation characterizes expressivity via density of realizable functions in spaces such as $C(K)$ for compact $K$. Recent theory refines the classical UAT by providing quantitative rates tied to architectural capacity and function-space regularity (e.g., Besov-type smoothness for ReLU networks) (Siegel, 2023), and by identifying minimal width conditions and sharp thresholds under which universality holds or fails even on compact domains (Kim et al., 2023). Extensions to broader settings consider uniform approximation beyond compact domains, where tail control becomes essential (Bournez et al., 2025), and alternative parameterizations that preserve approximation goals while altering the mechanism (Liu et al., 2024). Related ideas appear in operator learning, which elevates universality to maps between function spaces and connects approximation theory to PDE solution operators (Kovachki et al., 2023; Furuya et al., 2023, 2024b). Our work follows this modern viewpoint but changes the quantifier structure: rather than a weight-indexed family across tasks, we study the prompt-indexed hypothesis class induced by a fixed executor, asking when varying only the prompt yields density in $C(K)$ and what structural constraints govern such switching.
2.2 Expressivity and Approximation Results for Transformers
A growing theory treats Transformers as explicit function approximators and derives approximation guarantees under capacity and regularity assumptions (Takakura and Suzuki, 2023; Jiang and Li, 2024). Closely related are results on prompt- or context-conditioned families, where context acts as a control signal that expands the realizable set even with constrained core weights (Petrov et al., 2024; Furuya et al., 2024a). Complementary complexity- and mechanism-oriented analyses link Transformer expressivity to structured computation along the sequence and identify intrinsic limitations tied to width, depth, embedding dimension, and finite-precision or parallelizability constraints (Merrill and Sabharwal, 2023a; Sanford et al., 2023; Merrill and Sabharwal, 2023b). Our setting aligns with these perspectives but focuses on a single fixed executor with prompt-controlled computation, made explicit via a routing/arithmetic/composition decomposition.
2.3 Prompting as Adaptation: Soft Prompts, Prefix Tuning, and Prompt Tuning
Prompt-based adaptation is a widely used parameter-efficient alternative to full fine-tuning, and many methods can be viewed as updating a small task-specific parameterization while keeping most weights fixed (Ding et al., 2023). Recent variants design prompts to steer attention patterns (Wang et al., 2023), construct compositional or codebook-based prompts to control parameter growth (Lin et al., 2024), and stabilize prompt reliance to improve robustness and transfer (Li et al., 2024). Other work connects prompt tuning to low-rank reparameterizations and hybrid prompt modules that improve efficiency and flexibility (Jain et al., 2024; Tang et al., 2025), with analyses suggesting shared structure that benefits estimation and sample efficiency (Le et al., 2024). In contrast, we focus on expressivity: prompts index a hypothesis class under fixed weights, and we provide a constructive mechanism-level account of when prompt-only switching is possible in principle.
2.4 In-Context Learning and Prompting Theory
Theoretical work on in-context learning (ICL) interprets prompting as forward-pass adaptation and asks what algorithmic behavior is implemented when demonstrations are provided (Akyürek et al., 2022; Garg et al., 2022). Constructive and mechanistic analyses connect ICL to optimization dynamics and meta-learning views, showing that Transformer components can approximate gradient-based updates under suitable parameterizations (Von Oswald et al., 2023; Dai et al., 2023), and other perspectives emphasize compression into low-dimensional task representations (Hendel et al., 2023) or extend optimization views beyond linear settings (Cheng et al., 2023). Empirical and geometric probes further highlight robustness gaps and irregularities, motivating sharper conditions on what is learned in-context (Zhao et al., 2024). We are aligned with this agenda in treating prompting as computation, but we study an expressivity (existence) result for a fixed executor and make the switching mechanism explicit through routing/arithmetic/composition, rather than characterizing a particular learning rule or sample-efficiency regime.
2.5 Computational and Mechanistic Views: Attention as Retrieval and Transformers as Interpreters
Mechanistic-interpretability work increasingly models Transformers as decomposable computational systems, where attention behaves like structured routing and circuits implement recognizable subroutines on algorithmic tasks (Conmy et al., 2023; Nanda et al., 2023; Wang et al., 2022). In parallel, attention-as-retrieval abstractions connect factual recall and long-context behavior to controlled access to stored state or self-retrieval formalisms (Geva et al., 2023; Xu et al., 2023; Rubin and Berant, 2024), and causal analyses of induction heads provide fine-grained evidence for routing-centric interpretations of ICL (Crosbie and Shutova, 2025). Our lemma-based decomposition can be seen as a proof-level abstraction of these roles: prompt slots act as addressable memory, attention implements selective retrieval, the FFN performs local computation on retrieved fragments, and depth-wise composition accumulates these operations into a multi-step program.
3 Problem Setup
This section formalizes the approximation problem under the ”fixed-backbone + prompt” viewpoint. The key difference from the classical UAT, which changes weights across tasks, is that we ask whether a single fixed executor (backbone) can switch among different functions via prompts. To this end, we define the prompt-induced family of functions and take its density in $(C(\mathcal{X}),\|\cdot\|_{\infty})$ as our main objective.
3.1 Input domain and target function class
Assumption 3.1 (Input domain).
The input space is $\mathcal{X}\subset\mathbb{R}^{d}$, where $\mathcal{X}$ is compact.
Let $C(\mathcal{X})$ denote the space of real-valued continuous functions on $\mathcal{X}$, and measure approximation error in the sup-norm:
$\|f-g\|_{\infty}:=\sup_{x\in\mathcal{X}}|f(x)-g(x)|.$
Thus, for any $g\in C(\mathcal{X})$ and $\varepsilon>0$, the essence of the problem is whether the chosen model class can provide a function $f$ such that $\|f-g\|_{\infty}<\varepsilon$.
3.2 A fixed backbone and the prompt-induced hypothesis class
Definition 3.2 (Prompt space and prompt-conditional output).
Fix a maximum prompt length $L_{\max}\in\mathbb{N}$ and set $L:=L_{\max}$. Let the space of length-$L$ soft prompts (prefixes) be $\mathcal{P}:=\mathbb{R}^{L\times d_{p}}$. For tasks requiring fewer than $L_{\max}$ slots, we use zero-padding (null slots). For fixed parameters $\theta^{\ast}$, we denote the model’s scalar output by
$F_{\theta^{\ast}}:\mathcal{P}\times\mathcal{X}\to\mathbb{R},\qquad(p,x)\mapsto F_{\theta^{\ast}}(p,x).$
Assumption 3.3 (Well-definedness in $C(\mathcal{X})$).
For every prompt $p\in\mathcal{P}$, the map $x\mapsto F_{\theta^{\ast}}(p,x)$ is continuous on $\mathcal{X}$. Hence $\mathcal{H}_{\theta^{\ast}}\subset C(\mathcal{X})$ is well-defined.
Definition 3.4 (Prompt-induced hypothesis class).
We define the family of functions induced by a fixed backbone $\theta^{\ast}$ through prompts as
$\mathcal{H}_{\theta^{\ast}}:=\Big\{\,F_{\theta^{\ast}}(p,\cdot)\;:\;p\in\mathcal{P}\,\Big\}\subset C(\mathcal{X}).$
Definition 3.4 turns prompt engineering into a mathematical object. Prompt design can now be reduced to the question: ”how far can $\mathcal{H}_{\theta^{\ast}}$ reach?”
3.3 The Prompt-UAT objective
Remark 3.5 (Unconstrained prompts).
We allow the ambient prompt space $\mathcal{P}=\mathbb{R}^{L_{\max}\times d_{p}}$ without norm constraints to state a purely existential expressivity result; quantitative trade-offs (prompt norm/precision versus $\varepsilon$) are deferred. In the actual executor-level construction and error analysis, however, we work with an admissible subset $\mathcal{P}_{\mathrm{adm}}\subset\mathcal{P}$ produced by the encoding scheme, on which all stored records are uniformly bounded.
Clarification (why unconstrained prompts are not the punchline).
Remark 3.5 should be read as an upper-envelope expressivity statement: we first establish what is possible in principle when the prompt is treated as an external continuous control input. In particular, Theorem 5.1 does not imply that finite-context LLMs can realize arbitrary tasks; rather, it motivates asking how the best achievable error scales under realistic constraints (finite $L$, finite precision, norm bounds). Accordingly, in Section 7 and Section 8 we make explicit where costs enter (memory size via $L$, routing reliability via $\Delta/\tau$, and arithmetic approximation via FFN width), and defer the resulting trade-offs to follow-up work.
Definition 3.6 (Prompt-UAT property).
We say that a fixed backbone $\theta^{\ast}$ satisfies Prompt-UAT if
$\overline{\mathcal{H}_{\theta^{\ast}}}^{\|\cdot\|_{\infty}}=C(\mathcal{X}).$
Equivalently, for every $g\in C(\mathcal{X})$ and $\varepsilon>0$, there exists some $p_{g}\in\mathcal{P}$ such that $\|F_{\theta^{\ast}}(p_{g},\cdot)-g\|_{\infty}<\varepsilon$.
Remark 3.7 (Scalar output w.l.o.g.).
We state results for scalar outputs; vector-valued targets can be handled componentwise.
Remark 3.8 (Comparison to the classical UAT).
The classical UAT chooses parameters $\theta_{g}$ (depending on $g$) and approximates via $F_{\theta_{g}}$. In contrast, Prompt-UAT fixes a single $\theta^{\ast}$ and switches across targets $g$ using only the prompt $p$. That is, the weights fix the executor, while the prompt specifies the task instance.
Theorem 3.9 (Prompt-UAT (informal statement)).
For an appropriately designed (sufficiently large) Transformer family, there exists a fixed parameter $\theta^{\ast}$ such that Definition 3.6 holds. In particular, for every $g\in C(\mathcal{X})$ and every $\varepsilon>0$, there exists some $p_{g}\in\mathcal{P}$ satisfying
$\sup_{x\in\mathcal{X}}\big|F_{\theta^{\ast}}(p_{g},x)-g(x)\big|<\varepsilon.$
Remark 3.10 (Interpretation).
Theorem 3.9 formalizes the view that ”the prompt is a program and the fixed Transformer is an executor” via the density of $\mathcal{H}_{\theta^{\ast}}$ in $(C(\mathcal{X}),\|\cdot\|_{\infty})$. In subsequent sections, we provide a constructive proof by combining routing through attention, arithmetic through the FFN, and composition via block iteration.
4 Model
In this section we define a minimal Transformer variant that admits a direct constructive proof. Our block retains only (i) selective readout (routing) via self-attention, (ii) token-wise local computation (arithmetic/gating) via an FFN, and (iii) composition via depth-wise repetition. In subsequent sections, we isolate these three functionalities as separate lemmas, so that constructing the executor reduces to assembling them.
4.1 Architecture
Assumption 4.1 (Simplified Transformer).
We consider a stack of $T$ blocks. Each block consists of single-head self-attention and a token-wise two-layer ReLU FFN, and both modules include residual connections. We do not use LayerNorm or Dropout.
Why this simplification.
We remove LayerNorm/Dropout to isolate the minimal computational roles used in the construction: selective retrieval (attention), token-wise local computation (FFN), and depth-wise composition. These components already suffice to implement the interpreter blueprint of Section 7; architectural details can be reintroduced once the routing/arithmetic/composition core is fixed. Our goal here is not architectural realism but a clean decomposition that can serve as a proof-level scaffold.
Dimensions and tokens.
Let the input domain be $\mathcal{X}\subset\mathbb{R}^{d}$ (compact, as in Section 3), and let $D$ denote the model width. In this constructive section we take the prompt dimension to match the model width, i.e., $d_{p}=D$.
A prompt is a length-$L$ token sequence (matrix) $P=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$ with $p_{i}\in\mathbb{R}^{D}$. An input $x\in\mathcal{X}$ is embedded into a single input token $u(x)=E_{x}x+b_{x}\in\mathbb{R}^{D}$, where $E_{x}\in\mathbb{R}^{D\times d}$ and $b_{x}\in\mathbb{R}^{D}$. We also introduce an initial output token $o\in\mathbb{R}^{D}$ and a work token $w\in\mathbb{R}^{D}$ (initialized to $w=\mathbf{0}$), which serves as a scratch register for intermediate computations. Letting the sequence length be $n=L+3$, the initial state is
$Z^{(0)}(P,x)=\begin{bmatrix}p_{1}\\ \vdots\\ p_{L}\\ u(x)\\ w\\ o\end{bmatrix}\in\mathbb{R}^{n\times D}.$
Row-wise softmax with temperature.
For $\tau>0$, let $\mathcal{S}_{\tau}$ denote the row-wise softmax defined by $\mathcal{S}_{\tau}(r)=\mathrm{softmax}(r/\tau)$ for each row vector $r$, and applied row-wise to matrices.
Single-head attention.
Let $W_{Q},W_{K},W_{V}\in\mathbb{R}^{D\times D}$. Define single-head self-attention by
$\mathcal{A}_{\tau}(Z)$ $=\mathcal{S}_{1}\!\Bigg(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\tau\sqrt{D}}\Bigg)\,(ZW_{V})$
$=\mathcal{S}_{\tau}\!\Bigg(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\sqrt{D}}\Bigg)\,(ZW_{V}).$
Unless stated otherwise, we use full self-attention; in a causal variant, the output token can still attend to the prompt prefix.
Token-wise FFN.
For a hidden width $h$, define the token-wise two-layer ReLU FFN
$\phi(z)=W_{2}\sigma(W_{1}z+b_{1})+b_{2},\quad W_{1}\in\mathbb{R}^{h\times D},\;W_{2}\in\mathbb{R}^{D\times h},$
where $\sigma$ is applied elementwise. Extend $\phi$ row-wise to matrices by $(\Phi(Z))_{i,:}=\phi(Z_{i,:})$.
Block update and scalar readout.
The block update is written as
$Z^{(t+\frac{1}{2})}$ $=Z^{(t)}+\mathcal{A}_{\tau}\!\big(Z^{(t)}\big),$
$Z^{(t+1)}$ $=Z^{(t+\frac{1}{2})}+\Phi\!\big(Z^{(t+\frac{1}{2})}\big),\qquad t=0,\dots,T-1.$
From the final output token $z^{(T)}_{\mathrm{out}}:=Z^{(T)}_{n,:}$, we define the scalar output as
$F_{\theta}(p,x):=v_{\mathrm{out}}^{\top}z^{(T)}_{\mathrm{out}}+b_{\mathrm{out}}.$
The full parameter set $\theta$ includes all weights above as well as $(E_{x},b_{x},o,v_{\mathrm{out}},b_{\mathrm{out}})$.
4.2 Viewing the prompt as key–value memory
For the prompt matrix $P=[p_{1};\dots;p_{L}]\in\mathbb{R}^{L\times D}$, define slot-wise keys and values by
$K_{P}:=PW_{K},\qquad V_{P}:=PW_{V},$
where $W_{K},W_{V}$ are the same matrices used inside $\mathcal{A}_{\tau}$.
Given a query $q\in\mathbb{R}^{D}$, define the prompt read operation as
$\mathcal{R}_{\tau}(q;P):=\mathcal{S}_{\tau}\!\left(\frac{qK_{P}^{\top}}{\sqrt{D}}\right)\,V_{P}\in\mathbb{R}^{D}.$
Since $\mathcal{S}_{\tau}(\cdot)$ outputs a probability vector (nonnegative entries summing to $1$), $\mathcal{R}_{\tau}(q;P)$ is a convex combination of the rows of $V_{P}$. Hence, if a sufficient score margin is ensured, taking $\tau$ small yields near one-hot routing that concentrates on a single slot (Routing Lemma).
In our constructions, we partition the slots into parameter storage and work memory:
$L]=\mathcal{I}_{\mathrm{par}}\;\dot{\cup}\;\mathcal{I}_{\mathrm{wrk}}.$
In summary, the prompt stores a program (parameters/control signals), while attention retrieves the required pieces and feeds them into state updates.
5 Main Theorem
This section states the Prompt-UAT objective as a single theorem and summarizes the proof strategy within one page. The key point is that the weights $\theta^{\ast}$ are fixed as an executor, while task-specific degrees of freedom are supplied by the prompt $p$.
5.1 Theorem statement
We reuse the input domain $\mathcal{X}\subset\mathbb{R}^{d}$ (compact) and the prompt space $\mathcal{P}=\mathbb{R}^{L\times d_{p}}$ from Section 3. Assumption 3.3 ensures $\mathcal{H}_{\theta}\subset C(\mathcal{X})$ is well-defined.
Theorem 5.1 (Prompt-UAT (fixed backbone + prompt switching)).
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be compact and let $\mathcal{P}=\mathbb{R}^{L\times d_{p}}$. Within the simplified Transformer family of Section 4, there exist architecture sizes $(T,D,\text{FFN width})$ and a temperature $\tau>0$, and there exist fixed parameters $\theta^{\ast}$ (including $\tau$ if treated as a parameter) such that the following holds: for every $g\in C(\mathcal{X})$ and every $\varepsilon>0$, there exists a prompt $p_{g}\in\mathcal{P}$ satisfying
$\sup_{x\in\mathcal{X}}\big|F_{\theta^{\ast}}(p_{g},x)-g(x)\big|<\varepsilon.$
Equivalently, $\overline{\mathcal{H}_{\theta^{\ast}}}^{\|\cdot\|_{\infty}}=C(\mathcal{X})$.
Remark 5.2 (Expressivity versus learnability).
Theorem 5.1 is an expressivity (existence) result: it asserts that for each target $g$ there exists a prompt $p_{g}$ under a fixed executor $\theta^{\ast}$. It does not provide an algorithm to find $p_{g}$, nor does it claim that pretrained LLMs satisfy the theorem as stated. We view optimization and pretraining effects as separate axes that determine which parts of $\mathcal{H}_{\theta^{\ast}}$ are practically reachable.
5.2 Proof overview
The proof combines standard UAT with an interpreter construction.
Proposition 5.3 (Proof blueprint).
Fix $\theta^{\ast}$ as in Theorem 5.1. Then for any $g\in C(\mathcal{X})$ and any $\varepsilon>0$:
  1. 1.
    (MLP approximation) There exists a ReLU MLP $N_{g}$ such that $\|N_{g}-g\|_{\infty}<\varepsilon/2$.
  2. 2.
    (Executor approximation) There exists a prompt $p_{g}\in\mathcal{P}$ encoding the parameters of $N_{g}$ such that $\|F_{\theta^{\ast}}(p_{g},\cdot)-N_{g}\|_{\infty}<\varepsilon/2$.
  3. 3.
    (Combination) By the triangle inequality, $\|F_{\theta^{\ast}}(p_{g},\cdot)-g\|_{\infty}<\varepsilon$.
Proof sketch.
In the first step, we apply the standard universal approximation theorem for ReLU MLPs: for any $g\in C(\mathcal{X})$, choose an MLP $N_{g}$ satisfying $\|N_{g}-g\|_{\infty}<\varepsilon/2$. Next, we construct a fixed executor $\theta^{\ast}$. The executor interprets prompt slots as key–value memory, selectively retrieves the required parameter fragments via attention-based routing, uses the token-wise FFN as a local arithmetic module to approximate each layer’s linear transformations and ReLU gating, and composes these operations across blocks to unroll the forward pass of $N_{g}$. By allocating stepwise error budgets for routing and arithmetic, we can control the total accumulated error to be at most $\varepsilon/2$. Finally, applying the triangle inequality yields $\|F_{\theta^{\ast}}(p_{g},\cdot)-g\|_{\infty}<\varepsilon$. ∎
6 Key Lemmas: Routing, Arithmetic, and Composition
The executor construction for Theorem 5.1 can be summarized as a combination of three functionalities: (i) attention almost-selectively reads a desired prompt slot (routing), (ii) the FFN combines retrieved values with the current state to perform a local update (arithmetic), and (iii) when iterated, the accumulated error is controlled via an explicit budget (composition).
6.1 Routing lemma
We use $\pi_{\tau}$ for the softmax on a single score vector, and $S_{\tau}$ for the row-wise extension applied to matrices (Subsection 4.1); they coincide when restricted to one row.
Definition 6.1 (Softmax with temperature and margin).
For $\tau>0$ and $s\in\mathbb{R}^{L}$, define
$\pi_{\tau}(s)_{i}:=\frac{\exp(s_{i}/\tau)}{\sum_{j=1}^{L}\exp(s_{j}/\tau)}\,,\qquad i\in[L].$
For $i^{\star}\in[L]$, we say that $s$ has an $(i^{\star},\Delta)$-margin if $s_{i^{\star}}\geq\max_{j\neq i^{\star}}s_{j}+\Delta$.
Lemma 6.2 (Routing: exponential concentration).
If the $(i^{\star},\Delta)$-margin in Definition 6.1 holds, then
$1-\pi_{\tau}(s)_{i^{\star}}\leq(L-1)e^{-\Delta/\tau},\quad\pi_{\tau}(s)_{j}\leq e^{-\Delta/\tau}\ \ (j\neq i^{\star}).$
Hence, if $\Delta/\tau\to\infty$, then $\pi_{\tau}(s)\to e_{i^{\star}}$.
Definition 6.3 (Prompt retrieval).
Let the prompt matrix be $P\in\mathbb{R}^{L\times D}$ with slot-keys and slot-values
$K_{P}:=PW_{K},\qquad V_{P}:=PW_{V},$
and let $q\in\mathbb{R}^{D}$ be a query. Define the read operation by
$\mathcal{R}_{\tau}(q;P):=\pi_{\tau}(qK_{P}^{\top})\,V_{P}\in\mathbb{R}^{D}.$
Corollary 6.4 (Copying error).
Let $s(q;P):=qK_{P}^{\top}$. If $s(q;P)$ has an $(i^{\star},\Delta)$-margin, then for any norm $\|\cdot\|$,
$\big\|\mathcal{R}_{\tau}(q;P)-V_{P,i^{\star}}\big\|$
$\leq\big(1-\pi_{\tau}(s(q;P))_{i^{\star}}\big)\cdot\max_{j\in[L]}\big\|V_{P,j}-V_{P,i^{\star}}\big\|$
$\leq(L-1)e^{-\Delta/\tau}\cdot\max_{j\in[L]}\big\|V_{P,j}-V_{P,i^{\star}}\big\|.$
Remark 6.5 (Where the margin comes from).
In the constructive proof, the $(i^{\star},\Delta)$-margin is enforced by designing keys so that the intended slot has a strictly larger query–key score than all others (e.g., via nearly-orthogonal keys and aligned queries).
6.2 Arithmetic lemma
Lemma 6.6 (Approximation power of a 2-layer ReLU FFN).
Let $\mathcal{U}\subset\mathbb{R}^{d_{u}}$ be compact and let $T:\mathcal{U}\to\mathbb{R}^{D}$ be continuous. For any $\delta>0$, there exist a width $h$ and a 2-layer ReLU network $\Phi:\mathbb{R}^{d_{u}}\to\mathbb{R}^{D}$ such that
$\sup_{u\in\mathcal{U}}\|\Phi(u)-T(u)\|\leq\delta.$
Remark 6.7.
In our constructions, $\mathcal{U}$ is a compact set containing tuples of the form (retrieved parameter fragments, current state), and $T$ is the intended layerwise local update. Lemma 6.6 provides the required token-wise arithmetic module.
6.3 Composition lemma
Lemma 6.8 (Error accumulation under composition).
Let $\mathcal{Z}$ be a normed space and let $G_{t},\widetilde{G}_{t}:\mathcal{Z}\to\mathcal{Z}$. Define the compositions
$G_{1:T}:=G_{T}\circ\cdots\circ G_{1},\qquad\widetilde{G}_{1:T}:=\widetilde{G}_{T}\circ\cdots\circ\widetilde{G}_{1},$
and the stepwise errors $\epsilon_{t}:=\sup_{z\in\mathcal{Z}}\|\widetilde{G}_{t}(z)-G_{t}(z)\|$. If each $G_{t}$ is Lipschitz on $\mathcal{Z}$ with constant $L_{t}$, then
$\sup_{z\in\mathcal{Z}}\|\widetilde{G}_{1:T}(z)-G_{1:T}(z)\|\leq\sum_{t=1}^{T}\epsilon_{t}\prod_{s=t+1}^{T}L_{s}$
holds.
Remark 6.9 (Error budgeting).
In the executor construction, $\epsilon_{t}$ decomposes into a routing error (e.g., $O(e^{-\Delta/\tau})$) plus an arithmetic error (the FFN approximation accuracy). By allocating per-step budgets $\{\epsilon_{t}\}$ so that the bound in Lemma 6.8 is at most a target total error $\eta$, we control the overall executor error.
7 Construction: Prompt-Encoded MLP Execution
This section presents the skeleton of the interpreter (executor) construction. The key idea is to first approximate a target function $g$ by a ReLU MLP $N_{g}$, and then design a fixed Transformer $\theta^{\ast}$ that uniformly emulates the forward pass of $N_{g}$ by reading $N_{g}$’s parameters stored in the prompt. Throughout this section, the input domain is $\mathcal{X}\subset\mathbb{R}^{d}$ (compact) as in Section 3. We fix a maximum prompt length $L_{\max}$ (hence the available memory) and represent every prompt as a matrix in $\mathbb{R}^{L_{\max}\times D}$, using zero-padding in unused slots. Let $L\leq L_{\max}$ denote the number of nonzero (used) slots.
7.1 Encoding: Prompt as Parameter Memory
Definition 7.1 (Target MLP and boundedness).
Let $N:\mathbb{R}^{d}\to\mathbb{R}$ be a depth-$S$ ReLU MLP, where (by convention) $S$ counts the input layer, the (S2)(S\!-\!2) hidden ReLU layers, and the output layer. Let $m_{\ell}$ denote the width of hidden layer $\ell$ and set $m_{0}=d$. Define
$h^{(0)}(x)$ $=x,$
$h^{(\ell)}(x)$ $=\sigma\!\big(W^{(\ell)}h^{(\ell-1)}(x)+b^{(\ell)}\big),\quad\ell=1,\dots,S-2,$
$N(x)$ $=W^{(S-1)}h^{(S-2)}(x)+b^{(S-1)},$
where $\sigma=\mathrm{ReLU}$, $W^{(\ell)}\in\mathbb{R}^{m_{\ell}\times m_{\ell-1}}$ and $b^{(\ell)}\in\mathbb{R}^{m_{\ell}}$ for $\ell\leq S-2$, and the output layer satisfies $W^{(S-1)}\in\mathbb{R}^{1\times m_{S-2}}$, $b^{(S-1)}\in\mathbb{R}$. We assume that for some $\Lambda>0$ all parameters are uniformly bounded in the entrywise max norm: $\|W^{(\ell)}\|_{\max}\leq\Lambda$, $\|b^{(\ell)}\|_{\max}\leq\Lambda$ for all $\ell\in\{1,\dots,S-1\}$.
Running example (one hidden layer).
Consider a depth-$S=3$ ReLU MLP with one hidden layer:
$N(x)=a^{\top}\sigma(Wx+b)+c,$
which matches Definition 7.1 by identifying $W^{(1)}=W$, $b^{(1)}=b$, $W^{(2)}=a^{\top}$, and $b^{(2)}=c$. A concrete choice of chunking is to store (i) each row $W_{r:}$ and the corresponding bias entry $b_{r}$ as one fragment, (ii) the output weight $a_{r}$ as one fragment per hidden unit, and (iii) the scalar $c$ as a final fragment. Thus $\mathcal{J}$ can be indexed by hidden-unit IDs together with a small tag indicating whether the payload is $(W_{r:},b_{r})$ or $a_{r}$, plus one slot for $c$; the addressing map $\mathrm{addr}:\mathcal{J}\to[L_{\max}]$ places these fragments into predetermined prompt slots. During execution, the work token retrieves the fragments for unit $r$, computes $\langle W_{r:},x\rangle+b_{r}$, applies ReLU, multiplies by $a_{r}$, and accumulates across $r$; repeating this logic realizes the layerwise macro-step in Definition 7.7.
Definition 7.2 (Chunking and addressing).
Fix a finite index set $\mathcal{J}$ and a chunking map $j\mapsto\mathrm{payload}_{j}(N)\in\mathbb{R}^{q}$ that extracts parameter fragments of $N$ (e.g., blocks of weights/biases or individual rows/columns). Fix an addressing function $\mathrm{addr}:\mathcal{J}\to[L_{\max}]$ that assigns each fragment index $j$ to a prompt slot.
Lemma 7.3 (Encoding feasibility via coordinate splitting).
Let $D$ be the model width. Fix integers $d_{k},d_{v}$ with $D\geq d_{k}+d_{v}$. Choose $W_{K},W_{V}\in\mathbb{R}^{D\times D}$ so that they read disjoint coordinate blocks, e.g.,
$W_{K}=\begin{bmatrix}I_{d_{k}}&0\\ 0&0\end{bmatrix},\qquad W_{V}=\begin{bmatrix}0&0\\ 0&I_{d_{v}}\end{bmatrix}$
(after padding to $D\times D$). Let each prompt token be a row vector $p_{i}\in\mathbb{R}^{1\times D}$ and define
$p_{i}=\big[\ \kappa(i)^{\top}\ \ \nu(y_{i})^{\top}\ \ 0^{\top}\ \big]\in\mathbb{R}^{1\times D},$
where $\kappa(i)\in\mathbb{R}^{d_{k}}$ and $\nu(y_{i})\in\mathbb{R}^{d_{v}}$ are prescribed. Then $(p_{i}W_{K})_{1:d_{k}}=\kappa(i)$ and $(p_{i}W_{V})_{(d_{k}+1):(d_{k}+d_{v})}=\nu(y_{i})$ simultaneously.
Definition 7.4 (Encoding map).
Fix an addressing (key) scheme $\kappa:[L_{\max}]\to\mathbb{R}^{d_{k}}$ and a value embedding map $\nu:\mathbb{R}^{q}\to\mathbb{R}^{d_{v}}$. Assuming Lemma 7.3, define the encoding $p_{N}=\mathrm{Enc}(N)\in\mathbb{R}^{L_{\max}\times D}$ as follows: for each $j\in\mathcal{J}$, let $i=\mathrm{addr}(j)$ and set the value stored at slot $i$ to be $\nu(\mathrm{payload}_{j}(N))$, while the slot key is fixed as $\kappa(i)$.
Remark 7.5.
The addressing (key) scheme is fixed across tasks, and the task-specific freedom enters only through the stored values (payloads). Thus, a central design variable becomes when and which slots are read during execution.
7.2 Execution: Layerwise Emulation
Definition 7.6 (Registers via coordinate projections).
Assume the model width $D$ is large enough to decompose $\mathbb{R}^{D}$ into disjoint coordinate blocks
$\mathbb{R}^{D}\cong\mathbb{R}^{r_{0}}\oplus\mathbb{R}^{r_{1}}\oplus\cdots\oplus\mathbb{R}^{r_{M}}.$
For $\ell\in[M]$, define $\mathrm{Read}_{\ell}(z)$ as the projection onto block $\ell$, and define $\mathrm{Write}_{\ell}(z,v)$ as the operation that replaces block $\ell$ of $z$ by $v$ (keeping all other blocks unchanged).
Definition 7.7 (Layer-emulation macro-step).
Let the ideal layer update at depth $\ell$ be
$G_{\ell}(u)=\sigma\!\big(W^{(\ell)}u+b^{(\ell)}\big),$
where $u$ denotes the current hidden vector stored in designated registers of the work token. We denote by $\widetilde{G}_{\ell}$ the approximate map implemented by the Transformer by (i) retrieving the required parameter chunks from the prompt via routing and (ii) applying an FFN-based local update that depends on (retrieved fragments, current registers).
Proposition 7.8 (Interpreter stage).
Fix an MLP depth $S$ and parameter bound $\Lambda$. For any $\eta>0$, there exist model sizes $(T,D,\text{FFN width})$ and a temperature $\tau>0$, and there exist fixed parameters $\theta^{\ast}$ within the simplified Transformer family such that the following holds:
for every ReLU MLP $N$ satisfying Definition 7.1, there exists a prompt $p_{N}=\mathrm{Enc}(N)\in\mathbb{R}^{L_{\max}\times D}$ (using zero-padding if needed) such that
$\sup_{x\in\mathcal{X}}\big|F_{\theta^{\ast}}(p_{N},x)-N(x)\big|\leq\eta.$
Proof sketch.
We implement the forward pass of $N$ via macro-steps for layers $\ell=1,\dots,S-1$. At macro-step $\ell$, the work token forms queries so that, by the key design $\kappa$, the correct slot attains an $(i^{\star},\Delta)$-margin. Lemma 6.2 then yields exponentially accurate retrieval, and it suffices to pick $\tau$ so that $(L_{\max}-1)e^{-\Delta/\tau}\leq\eta_{\mathrm{route}}$.
On the resulting compact set of register states and retrieved fragments (from bounded parameters and compact $\mathcal{X}$), Lemma 6.6 gives a token-wise FFN approximating the intended local update map $G_{\ell}$ within the arithmetic budget. Finally, Lemma 6.8 allocates per-step budgets so that routing and arithmetic errors accumulate to at most $\eta$, yielding $\sup_{x\in\mathcal{X}}\big|F_{\theta^{\ast}}(p_{N},x)-N(x)\big|\leq\eta$. ∎
Remark 7.9 (Putting errors together).
For any $g\in C(\mathcal{X})$, let $N_{g}$ satisfy $\|N_{g}-g\|_{\infty}<\varepsilon/2$ by the standard UAT for ReLU MLPs. Applying Proposition 7.8 with $\eta=\varepsilon/2$ gives
$\|F_{\theta^{\ast}}(p_{N_{g}},\cdot)-g\|_{\infty}$ $\leq\|F_{\theta^{\ast}}(p_{N_{g}},\cdot)-N_{g}\|_{\infty}$
$\quad+\|N_{g}-g\|_{\infty}$
$<\varepsilon.$
which implies Theorem 5.1.
8 Discussion and Limitations
Discussion.
We position prompt engineering as a theoretical object: for a fixed backbone, varying the prompt induces a well-defined hypothesis class. This reframes the prompt as an external parameter—a program input—so expressivity questions can be posed directly about a single mathematical object.
Our contribution is constructive and mechanism-oriented. The proof isolates a reusable decomposition of prompt-based switching: prompt slots serve as addressable memory, attention performs routing via near-selective retrieval, the FFN performs arithmetic via local computation conditioned on retrieved fragments, and depth implements composition by accumulating these updates into multi-stage computation. Consequently, the construction makes explicit what is stored, how it is selected, and how it is transformed across layers.
This viewpoint also clarifies where practical constraints enter—memory via $L$, routing reliability via $\Delta/\tau$, and arithmetic accuracy via FFN width—even though we do not optimize these costs here. By turning informal prompt ”tips” into explicit knobs, it sets up subsequent work to quantify scaling laws and lower bounds under realistic constraints.
Limitations.
Our theorem is existential and concerns expressivity; it should not be read as an empirical claim that pretrained LLMs satisfy Prompt-UAT as stated. Moreover, while we prove that an appropriate prompt exists for each target, we do not provide efficient procedures for finding such prompts. We also do not quantify the costs induced by realistic constraints—such as prompt length, finite precision, norm bounds, or the margins/temperature scaling required for reliable routing—within the main text.
Future Work.
A direct next step is to quantify trade-offs between prompt length/precision/norm constraints and approximation error $\varepsilon$. A second direction is to derive lower bounds and impossibility results for hard prompts (finite vocabularies and finite lengths), including information-theoretic limits on how much task specification can be conveyed. A third direction is to connect existence to learnability: whether the existential prompts can be recovered by practical optimization, and how pretrained constraints (architecture, training-induced biases, and optimization dynamics) alter the boundary between possible and impossible switching.
9 Conclusion
We introduced a framework that treats prompt engineering as a coherent theoretical object rather than an empirical craft. Instead of changing weights across tasks, we ask what functionality can be switched when a single backbone is kept fixed as an executor and the prompt is used as an external program. We formalize the family of behaviors obtained by varying prompts under a fixed backbone and prove a constructive existential guarantee: by changing only the prompt, the fixed executor can approximate a broad class of targets to arbitrary accuracy. The construction makes the switching mechanism explicit—prompt slots act as addressable memory, attention provides selective routing, the FFN provides local arithmetic, and depth-wise repetition composes these updates into a full computation. This viewpoint motivates future work on trade-offs under length, precision, and norm constraints, limits for finite hard prompts, pretraining-induced restrictions, and the learnability of effective prompts under practical optimization.
Impact Statement
We study prompt engineering as an expressivity problem: what a fixed backbone can realize when only the prompt is varied. Our impact is primarily conceptual and theoretical. By formalizing the prompt-induced hypothesis class and proving a constructive existence result for prompt-based switching, we provide precise language for reasoning about what prompts can enable or restrict in principle, and for separating expressivity questions from optimization or training effects.
This work is foundational rather than deployment-facing. We do not claim that pretrained large language models satisfy the stated property, and we do not provide procedures for finding effective prompts under realistic constraints such as length, precision, or norm bounds. Accordingly, the results should not be interpreted as evidence of practical capability, nor as any form of safety, robustness, or reliability guarantee for prompt-only control.
Potential downstream benefits include improved interpretability of prompt-conditioned behavior, clearer theoretical guidance for prompt tuning and prompt search, and more principled characterization of failure modes and invariants in prompt-controlled systems. Potential risks include misreading the existence result as a practical capability claim or using it to justify overconfident reliance on prompt-only control in safety-critical settings. We mitigate these risks by clearly delimiting the scope of the theorem, emphasizing the absence of constraint-aware and learnable guarantees, and explicitly identifying robustness, constrained trade-offs, and learnability as open directions.
Appendix A Notation and Basic Operators
This appendix fixes notation and frequently used operators. Its role is to centralize symbols and conventions so that the subsequent appendices can state routing, computation, and composition claims without repeatedly reintroducing definitions. We work with a compact input domain $X\subset\mathbb{R}^{d}$ and the prompt-conditioned map $F_{\theta^{\ast}}$ with hypothesis class $H_{\theta^{\ast}}$ as introduced in the main text.
Basic symbols.
Scalars are denoted by lowercase letters (e.g., $\varepsilon,\tau,\Delta$). Vectors are bold lowercase (e.g., $\mathbf{q}$) and matrices are bold uppercase (e.g., $\mathbf{W}$). For a vector $\mathbf{u}$, $\mathbf{u}_{i}$ denotes its $i$-th coordinate; for a matrix $\mathbf{M}$, $\mathbf{M}_{ij}$ denotes its $(i,j)$ entry.
Domain, function space, and norms.
Inputs are $x\in X\subset\mathbb{R}^{d}$ (compact; see Assumption 3.1), and the target function is $g\in C(X)$. For $f:X\to\mathbb{R}$, the sup norm is
$\|f\|_{\infty}\;:=\;\sup_{x\in X}|f(x)|.$
For vectors, $\|\cdot\|_{2}$ and $\|\cdot\|_{\infty}$ denote the Euclidean and $\ell_{\infty}$ norms, respectively.
Fixed backbone and prompt-conditioning.
The fixed executor parameters are $\theta^{\ast}$. Prompts are soft prefixes of length-$L$ in the space $\mathcal{P}=\mathbb{R}^{L\times d_{p}}$ as in Definition 3.2, and we write $p=(p_{1},\dots,p_{L})\in\mathcal{P}$. The prompt-conditioned scalar output is $F_{\theta^{\ast}}(p,x)$, and the induced prompt-switching family is $H_{\theta^{\ast}}=\{F_{\theta^{\ast}}(p,\cdot):p\in P\}\subset C(X)$ as in Definition 3.4.
Error budgets.
We use $\varepsilon>0$ for an end-to-end uniform error target. When splitting the budget, we write $\varepsilon_{\mathrm{approx}}$ for the approximation error (e.g., $g\approx N_{g}$) and $\varepsilon_{\mathrm{exec}}$ for the executor emulation error, with $\varepsilon_{\mathrm{approx}}+\varepsilon_{\mathrm{exec}}\leq\varepsilon$. Local budgets are denoted by $\delta$ or $\eta$.
Row-wise softmax and attention operators.
Let $S_{\tau}$ denote the row-wise softmax with temperature $\tau>0$, i.e., $S_{\tau}(r)=\mathrm{softmax}(r/\tau)$ applied to each row. Let the model width be $D$, and let $W_{Q},W_{K},W_{V}\in\mathbb{R}^{D\times D}$. For a token/state matrix $Z\in\mathbb{R}^{n\times D}$, define single-head self-attention by
$A_{\tau}(Z)\;:=\;S_{\tau}\!\left(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\sqrt{D}}\right)\,(ZW_{V}).$
Prompt read as memory access.
When prompts are treated as tokens in $\mathbb{R}^{D}$ (typically by taking $d_{p}=D$ or by an implicit linear embedding), we write the prompt matrix as $\textbf{P}=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$ and define
$K_{\textbf{P}}:=\textbf{P}W_{K},\qquad V_{\textbf{P}}:=\textbf{P}W_{V}.$
Given a query $q\in\mathbb{R}^{D}$, the prompt read operation is
$\mathcal{R}_{\tau}(q;\textbf{P})\;:=\;S_{\tau}(qK_{P}^{\top})\,V_{P}\in\mathbb{R}^{D}.$
Routing analyses use a score margin $\Delta>0$ and an impurity level $\rho\in(0,1)$ ($1-\rho$ nearly one-hot”).
ReLU and FFN.
The ReLU is $\sigma(u)=\max\{u,0\}$. A token-wise two-layer ReLU FFN with hidden width $h$ is
$\phi(z)\;:=\;W_{2}\,\sigma(W_{1}z+b_{1})+b_{2},\qquad W_{1}\in\mathbb{R}^{h\times D},\;W_{2}\in\mathbb{R}^{D\times h},$
applied row-wise to token matrices (by applying $\phi$ to each row).
A.1 Symbol table
Table 1: Core notation used throughout the paper.
Symbol Meaning
$X\subset\mathbb{R}^{d}$ Compact input domain.
$d$ Input dimension.
$x\in X$ Input point.
$C(X)$ Continuous real-valued functions on $X$.
$\|f\|_{\infty}$ Sup norm $\sup_{x\in X}|f(x)|$.
$\theta^{\ast}$ Fixed backbone/executor parameters.
$\mathcal{P}=\mathbb{R}^{L\times d_{p}}$ Soft-prompt space; $L$ prompt length, $d_{p}$ prompt width.
$p=(p_{1},\ldots,p_{L})\in P$ A prompt (soft prefix).
$F_{\theta^{\ast}}(p,x)$ Prompt-conditioned scalar output.
$H_{\theta^{\ast}}$ Prompt-induced family $\{F_{\theta^{\ast}}(p,\cdot):p\in P\}\subset C(X)$.
$g\in C(X)$ Target function.
$N_{g}$ ReLU MLP approximator of $g$ (classical UAT step).
$\varepsilon$ End-to-end uniform error target.
$\varepsilon_{\mathrm{approx}},\varepsilon_{\mathrm{exec}}$ Approximation versus executor error budgets.
$\delta,\eta$ Local/intermediate error budgets.
$D$ Model width (token/state dimension).
$Z\in\mathbb{R}^{n\times D}$ Token/state matrix; $n$ is number of tokens.
$S_{\tau}$ Row-wise softmax with temperature $\tau>0$.
$\tau$ Temperature parameter (smaller $\tau$ sharpens routing).
$W_{Q},W_{K},W_{V}$ Attention projection matrices.
$A_{\tau}(Z)$ Single-head self-attention operator.
$\textbf{P}=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$ Prompt viewed in model space (when $d_{p}=D$ or embedded).
$K_{\textbf{P}}=\textbf{P}W_{K},\;V_{\textbf{P}}=\textbf{P}W_{V}$ Prompt keys/values.
$q\in\mathbb{R}^{D}$ Query vector.
$\mathcal{R}_{\tau}(q;\textbf{P})$ Prompt read $S_{\tau}(qK_{P}^{\top})V_{P}$.
$\Delta,\rho$ Routing margin and impurity level.
$\sigma(\cdot)$ ReLU activation.
$\phi(\cdot)$ Token-wise two-layer ReLU FFN; $h$ is hidden width.
Appendix B Simplified Transformer Architecture
This appendix specifies the simplified Transformer family used throughout the construction. The purpose is to pin down the minimal architectural components needed by the executor: row-wise softmax $S_{\tau}$, single-head self-attention $A_{\tau}$, the prompt read operation $\mathcal{R}_{\tau}(\cdot;P)$, and a token-wise two-layer ReLU FFN $\phi(\cdot)$ (extended row-wise as $\Phi(\cdot)$), all at a fixed model width $D$.
Tokens, prompt, and input embedding.
We take the prompt width to match the model width, i.e., $d_{p}=D$, so a prompt is a length-$L$ token sequence $p=(p_{1},\ldots,p_{L})\in\mathcal{P}=\mathbb{R}^{L\times D}$ with $p_{i}\in\mathbb{R}^{D}$. An input $x\in\mathcal{X}\subset\mathbb{R}^{d}$ is embedded into a single input token
$u(x)\;:=\;E_{x}x+b_{x}\in\mathbb{R}^{D},\qquad E_{x}\in\mathbb{R}^{D\times d},\ \ b_{x}\in\mathbb{R}^{D},$
and we introduce an initial output token $o\in\mathbb{R}^{D}$. Let the total sequence length be $n:=L+3$. The initial state (token matrix) is
$Z^{(0)}(p,x)\;:=\;\begin{bmatrix}p_{1}\\ \vdots\\ p_{L}\\ u(x)\\ w\\ o\end{bmatrix}\in\mathbb{R}^{n\times D}.$
We denote the $i$-th token (row) at depth $t$ by $z^{(t)}_{i}\in\mathbb{R}^{D}$, i.e., $Z^{(t)}_{i,:}=z^{(t)}_{i}$.
Row-wise softmax with temperature.
For $\tau>0$, $S_{\tau}$ denotes the row-wise softmax defined by $S_{\tau}(r)=\mathrm{softmax}(r/\tau)$ for each row vector $r$, applied row-wise to matrices.
Single-head self-attention.
Let $W_{Q},W_{K},W_{V}\in\mathbb{R}^{D\times D}$. Define single-head self-attention by
$A_{\tau}(Z)\;:=\;S_{\tau}\!\left(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\sqrt{D}}\right)\,(ZW_{V}),\qquad Z\in\mathbb{R}^{n\times D}.$
(Equivalently, the $\sqrt{D}$ scaling can be absorbed into $\tau$; we keep it explicit for clarity.)
Token-wise FFN.
For a hidden width $h$, define the token-wise two-layer ReLU FFN
$\phi(z)\;:=\;W_{2}\,\sigma(W_{1}z+b_{1})+b_{2},\qquad W_{1}\in\mathbb{R}^{h\times D},\;W_{2}\in\mathbb{R}^{D\times h},$
where $\sigma$ is applied elementwise. Extend $\phi$ row-wise to matrices by
$(\Phi(Z))_{i,:}\;:=\;\phi(Z_{i,:})\qquad(i=1,\ldots,n).$
Block update and scalar readout.
We consider a stack of $T$ blocks. The block update is written as
$Z^{(t+\frac{1}{2})}\;=\;Z^{(t)}+A_{\tau}(Z^{(t)}),\qquad Z^{(t+1)}\;=\;Z^{(t+\frac{1}{2})}+\Phi(Z^{(t+\frac{1}{2})}),\qquad t=0,\ldots,T-1.$
From the final output token $z^{(T)}_{\mathrm{out}}:=Z^{(T)}_{n,:}$, we define the scalar output as
$F_{\theta^{\ast}}(p,x)\;:=\;v_{\mathrm{out}}^{\top}z^{(T)}_{\mathrm{out}}+b_{\mathrm{out}},$
where $v_{\mathrm{out}}\in\mathbb{R}^{D}$ and $b_{\mathrm{out}}\in\mathbb{R}$ are fixed and included in $\theta^{\ast}$.
Viewing the prompt as key–value memory and prompt read.
For the prompt matrix $P=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$, define slot-wise keys and values by
$K_{P}\;:=\;PW_{K},\qquad V_{P}\;:=\;PW_{V},$
where $W_{K},W_{V}$ are the same matrices used inside $A_{\tau}$. Given a query $q\in\mathbb{R}^{D}$, define the prompt read operation
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}\!\left(\frac{qK_{P}^{\top}}{\sqrt{D}}\right)\,V_{P}\in\mathbb{R}^{D}.$
(As in self-attention, the $\sqrt{D}$ factor can be absorbed into $\tau$; we keep it explicit for consistency with $A_{\tau}$.)
Remark (simplifications).
We do not use LayerNorm, dropout, or multi-head attention. The construction uses only (i) selective readout via attention, (ii) token-wise local computation via the FFN, and (iii) depth-wise composition via residual stacking.
Appendix C Routing Lemma: Prompt-Memory Retrieval
This appendix formalizes the routing primitive used throughout the construction. The key object is the prompt read operation
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}\!\left(\frac{qK_{P}^{\top}}{\sqrt{D}}\right)\,V_{P}\in\mathbb{R}^{D},$
where $S_{\tau}$ is the row-wise softmax with temperature $\tau>0$ and where, for a prompt matrix $P=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$, we define slot-wise keys and values by
$K_{P}:=PW_{K},\qquad V_{P}:=PW_{V},$
using the same matrices $W_{K},W_{V}\in\mathbb{R}^{D\times D}$ as in $\mathcal{A}_{\tau}$.
C.1 Margin condition and impurity
Fix a query $q\in\mathbb{R}^{D}$ and define the score of slot $j\in[L]$ by
$s_{j}\;:=\;\langle q,(K_{P})_{j,:}\rangle\;=\;(qK_{P}^{\top})_{j}.$
Let $j^{\ast}\in[L]$ denote the intended slot. We say that $(q,P)$ satisfies a margin $\Delta>0$ for $j^{\ast}$ if
[…]
Let $\alpha:=S_{\tau}(qK_{P}^{\top})\in\mathbb{R}^{1\times L}$ be the attention weights over prompt slots; then
$\mathcal{R}_{\tau}(q;P)\;=\;\sum_{j=1}^{L}\alpha_{j}(V_{P})_{j,:}$
is a convex combination of prompt values. We refer to $1-\alpha_{j^{\ast}}$ as the routing impurity.
C.2 Routing bounds
Lemma C.1 (Softmax routing under a margin).
Fix $\tau>0$ and let $\alpha=S_{\tau}(qK_{P}^{\top})$. If the margin condition (1) holds for some $\Delta>0$, then
$\alpha_{j^{\ast}}$ $\;\geq\;\frac{1}{1+(L-1)e^{-\Delta/\tau}},$ (2)
$1-\alpha_{j^{\ast}}$ $\;\leq\;\frac{(L-1)e^{-\Delta/\tau}}{1+(L-1)e^{-\Delta/\tau}}\;\leq\;(L-1)e^{-\Delta/\tau}.$ (3)
Proof.
Let $M:=\max_{j\neq j^{\ast}}s_{j}$. By (1), $s_{j^{\ast}}\geq M+\Delta$. Hence,
$\sum_{j\neq j^{\ast}}e^{s_{j}/\tau}\;\leq\;(L-1)e^{M/\tau}\;\leq\;(L-1)e^{(s_{j^{\ast}}-\Delta)/\tau}\;=\;(L-1)e^{s_{j^{\ast}}/\tau}e^{-\Delta/\tau}.$
Therefore,
$\alpha_{j^{\ast}}=\frac{e^{s_{j^{\ast}}/\tau}}{e^{s_{j^{\ast}}/\tau}+\sum_{j\neq j^{\ast}}e^{s_{j}/\tau}}\;\geq\;\frac{1}{1+(L-1)e^{-\Delta/\tau}},$
and the impurity bound follows by $1-\alpha_{j^{\ast}}$. ∎
Lemma C.2 (Readout error under bounded values).
Assume the prompt values are uniformly bounded as $\|(V_{P})_{j,:}\|_{\infty}\leq B$ for all $j\in[L]$. Let $j^{\ast}$ satisfy (1), and let $\alpha=S_{\tau}(qK_{P}^{\top})$. Then
$\bigl\|\mathcal{R}_{\tau}(q;P)-(V_{P})_{j^{\ast},:}\bigr\|_{\infty}\;\leq\;2B\,(1-\alpha_{j^{\ast}})\;\leq\;2B\,(L-1)e^{-\Delta/\tau}.$ (4)
Proof.
Write $v_{j}:=(V_{P})_{j,:}\in\mathbb{R}^{D}$ and $v^{\ast}:=v_{j^{\ast}}$. Then
$\mathcal{R}_{\tau}(q;P)-v^{\ast}=(\alpha_{j^{\ast}}-1)v^{\ast}+\sum_{j\neq j^{\ast}}\alpha_{j}v_{j}.$
Taking $\|\cdot\|_{\infty}$ and using $\|v_{j}\|_{\infty}\leq B$ gives
$\|\mathcal{R}_{\tau}(q;P)-v^{\ast}\|_{\infty}\leq(1-\alpha_{j^{\ast}})\|v^{\ast}\|_{\infty}+\sum_{j\neq j^{\ast}}\alpha_{j}\|v_{j}\|_{\infty}\leq 2B(1-\alpha_{j^{\ast}}),$
and the exponential bound follows from (3). ∎
C.3 A convenient parameter choice
It is often convenient to target a desired impurity level $\rho\in(0,1)$.
Corollary C.3 (Temperature for a desired impurity).
Under the margin condition (1), if
$\tau\;\leq\;\frac{\Delta}{\log\!\bigl(\frac{L-1}{\rho}\bigr)},$
then $1-\alpha_{j^{\ast}}\leq\rho$. Consequently, under the bounded-value assumption of Lemma C.2,
$\bigl\|\mathcal{R}_{\tau}(q;P)-(V_{P})_{j^{\ast},:}\bigr\|_{\infty}\;\leq\;2B\rho.$
Proof.
By Lemma C.1, it suffices to ensure $(L-1)e^{-\Delta/\tau}\leq\rho$, which is equivalent to the stated condition on $\tau$. The readout bound then follows from Lemma C.2. ∎
Remark (enforcing the margin).
In the construction, prompts store keys such that the query $q$ produced at a given step aligns with the intended key and is separated by a margin $\Delta$ from all distractors. The lemmas above isolate the analytic consequence: routing error decays exponentially in $\Delta/\tau$.
Appendix D Arithmetic Lemma: Token-wise FFN Computation
This appendix formalizes the arithmetic primitive used in the construction. After routing injects the required information into a token, a token-wise two-layer ReLU FFN can approximate the desired local update uniformly on a bounded domain. Concretely, we treat the FFN as a map
$\phi(z)\;:=\;W_{2}\,\sigma(W_{1}z+b_{1})+b_{2},\qquad z\in\mathbb{R}^{D},$
and show that, for any continuous update rule on a bounded state set, one can choose $(W_{1},W_{2},b_{1},b_{2})$ so that $\phi$ matches that update to any prescribed uniform accuracy.
D.1 Bounded domain
For $B>0$, define the compact box
$\mathcal{B}_{B}\;:=\;\{z\in\mathbb{R}^{D}:\ \|z\|_{\infty}\leq B\}=[-B,B]^{D}.$
All uniform approximation statements below are taken over $\mathcal{B}_{B}$.
D.2 Piecewise-linear approximation
We first recall that continuous functions on a compact box admit uniform approximation by piecewise-linear maps.
Lemma D.1 (PL approximation on a box).
Let $f:\mathcal{B}_{B}\to\mathbb{R}$ be continuous. For any $\delta>0$, there exists a continuous piecewise-linear function $\tilde{f}:\mathcal{B}_{B}\to\mathbb{R}$ such that
$\sup_{z\in\mathcal{B}_{B}}\ |f(z)-\tilde{f}(z)|\;\leq\;\delta.$
Proof.
Since $f$ is continuous on the compact set $\mathcal{B}_{B}$, it is uniformly continuous. Choose $\eta>0$ such that $\|z-z^{\prime}\|_{\infty}\leq\eta$ implies $|f(z)-f(z^{\prime})|\leq\delta/2$.
Partition each coordinate interval [B,B][-B,B] into subintervals of length at most $\eta$ and form the induced rectangular grid on $\mathcal{B}_{B}$. Triangulate each grid cell into finitely many simplices (e.g., by ordering coordinates). Define $\tilde{f}$ by setting $\tilde{f}$ to equal $f$ at every grid vertex and extending $\tilde{f}$ linearly over each simplex.
Then $\tilde{f}$ is continuous and piecewise-linear. For any $z\in\mathcal{B}_{B}$, let $v$ be a vertex of the simplex containing $z$. By construction, $\|z-v\|_{\infty}\leq\eta$, hence
$|f(z)-f(v)|\leq\delta/2.$
Moreover, $\tilde{f}(v)=f(v)$ for all vertices $v$. Since $\tilde{f}$ is an affine combination of its vertex values on each simplex, there exists some vertex $v$ of that simplex such that
$|\tilde{f}(z)-\tilde{f}(v)|\leq\max_{v^{\prime}\ \text{vertex of simplex}}|\tilde{f}(v^{\prime})-\tilde{f}(v)|=\max_{v^{\prime}\ \text{vertex}}|f(v^{\prime})-f(v)|.$
All vertices of a simplex lie within $\ell_{\infty}$-distance at most $\eta$ from each other, so by uniform continuity $|f(v^{\prime})-f(v)|\leq\delta/2$. Therefore $|\tilde{f}(z)-f(v)|\leq\delta/2$ and
$|f(z)-\tilde{f}(z)|\leq|f(z)-f(v)|+|f(v)-\tilde{f}(z)|\leq\delta/2+\delta/2=\delta.$
D.3 ReLU representation of piecewise-linear functions
Next we use that two-layer ReLU networks represent continuous piecewise-linear functions.
Lemma D.2 (Two-layer ReLU realizes PL functions).
Let $\tilde{f}:\mathbb{R}^{D}\to\mathbb{R}$ be continuous and piecewise-linear with finitely many linear regions. Then there exist $m\in\mathbb{N}$, $w_{1},\ldots,w_{m}\in\mathbb{R}^{D}$, $a_{1},\ldots,a_{m}\in\mathbb{R}$, and an affine map $\ell(z)=\langle c,z\rangle+d$ such that
$\tilde{f}(z)\;=\;\ell(z)+\sum_{r=1}^{m}a_{r}\,\sigma(\langle w_{r},z\rangle)\qquad\text{for all }z\in\mathbb{R}^{D}.$
Equivalently, $\tilde{f}$ can be written as $W_{2}\sigma(W_{1}z+b_{1})+b_{2}$ for suitable matrices/vectors.
Proof.
A standard representation result for continuous piecewise-linear functions states that any such $\tilde{f}$ admits a ”hinge” decomposition into an affine function plus a finite sum of signed ReLU ridge functions. One way to see this is to observe that the set of break hyperplanes (where the gradient changes) is finite; across each such hyperplane the function has an affine kink. Each kink can be expressed by a term of the form $a\,\sigma(\langle w,z\rangle)$ (possibly after subtracting its affine part), and summing all kinks together with a base affine function yields $\tilde{f}$ exactly.
For completeness, we give an explicit reduction to the two-layer form. Write the above representation as
$\tilde{f}(z)=\langle c,z\rangle+d+\sum_{r=1}^{m}a_{r}\,\sigma(\langle w_{r},z\rangle).$
Let
$W_{1}:=\begin{bmatrix}w_{1}^{\top}\\ \vdots\\ w_{m}^{\top}\end{bmatrix}\in\mathbb{R}^{m\times D},\qquad b_{1}:=0\in\mathbb{R}^{m},\qquad W_{2}:=\begin{bmatrix}a_{1}&\cdots&a_{m}\end{bmatrix}\in\mathbb{R}^{1\times m},$
and set $b_{2}:=d$ and incorporate the affine term $\langle c,z\rangle$ by using the identity $\langle c,z\rangle=\langle c,\sigma(z)\rangle-\langle c,\sigma(-z)\rangle$, which is itself a two-layer ReLU network (Lemma D.4 below). Combining these gives a two-layer ReLU realization of $\tilde{f}$. ∎
D.4 Main arithmetic approximation lemma (vector-valued)
We now combine the previous lemmas and extend coordinate-wise to $\mathbb{R}^{D}$ outputs.
Lemma D.3 (Token-wise ReLU FFN approximates continuous updates).
Fix $B>0$. Let $f:\mathcal{B}_{B}\to\mathbb{R}^{D}$ be continuous and write $f=(f_{1},\ldots,f_{D})$. For any $\delta>0$, there exists a two-layer ReLU network $\phi:\mathbb{R}^{D}\to\mathbb{R}^{D}$ of the form $\phi(z)=W_{2}\sigma(W_{1}z+b_{1})+b_{2}$ such that
$\sup_{z\in\mathcal{B}_{B}}\ \|\phi(z)-f(z)\|_{\infty}\;\leq\;\delta.$
Proof.
Let $\delta>0$ be given. For each coordinate $i\in[D]$, apply Lemma D.1 to $f_{i}$ with accuracy $\delta$ to obtain a continuous piecewise-linear $\tilde{f}_{i}$ with
$\sup_{z\in\mathcal{B}_{B}}|f_{i}(z)-\tilde{f}_{i}(z)|\leq\delta.$
By Lemma D.2, each $\tilde{f}_{i}$ can be realized exactly by a two-layer ReLU network $\phi_{i}:\mathbb{R}^{D}\to\mathbb{R}$.
Define $\phi:\mathbb{R}^{D}\to\mathbb{R}^{D}$ by stacking the coordinate networks:
$\phi(z):=\bigl(\phi_{1}(z),\ldots,\phi_{D}(z)\bigr).$
This is again a two-layer ReLU network (block-diagonal construction of $W_{2}$), and for every $z\in\mathcal{B}_{B}$,
$\|\phi(z)-f(z)\|_{\infty}=\max_{i\in[D]}|\phi_{i}(z)-f_{i}(z)|=\max_{i\in[D]}|\tilde{f}_{i}(z)-f_{i}(z)|\leq\delta.$
D.5 Useful explicit gadget: exact affine maps
We also record a simple exact representation that is occasionally convenient.
Lemma D.4 (Exact affine maps).
For any matrix $U\in\mathbb{R}^{D\times D}$ and vector $c\in\mathbb{R}^{D}$, there exists a two-layer ReLU network $\psi(z)=\widetilde{W}_{2}\sigma(\widetilde{W}_{1}z+\widetilde{b}_{1})+\widetilde{b}_{2}$ such that $\psi(z)=Uz+c$ for all $z\in\mathbb{R}^{D}$.
Proof.
Use $z=\sigma(z)-\sigma(-z)$ coordinate-wise. Let
$\widetilde{W}_{1}=\begin{bmatrix}I\\ -I\end{bmatrix}\in\mathbb{R}^{2D\times D},\quad\widetilde{b}_{1}=0,\quad\widetilde{W}_{2}=\begin{bmatrix}U&-U\end{bmatrix}\in\mathbb{R}^{D\times 2D},\quad\widetilde{b}_{2}=c.$
Then $\widetilde{W}_{2}\sigma(\widetilde{W}_{1}z)=U\sigma(z)-U\sigma(-z)=Uz$, hence $\psi(z)=Uz+c$. ∎
D.6 How this lemma is used
In the executor, routing contributes a read vector $\mathcal{R}_{\tau}(q;P)$ to a designated work token via residual addition (see Appendix B). Thus the input to the FFN at that token lies in a bounded box $\mathcal{B}_{B}$ by design, encoding both the current work state and the freshly read parameters. Lemma D.3 then guarantees that a fixed token-wise FFN can approximate the desired local update uniformly to within $\delta$, which is the per-step arithmetic error used in the composition/error-budget argument.
Appendix E Composition and Error Budget Template
This appendix provides a reusable template for composing routing and token-wise computation while controlling the end-to-end error. The purpose is not to optimize constants, but to provide a clear bookkeeping scheme for allocating local error budgets across macro-steps and turning per-step guarantees into a uniform end-to-end bound.
E.1 Abstract macro-step model
Fix a prompt $p\in P$ and an input $x\in X$. Consider a designated work token whose state after macro-step $t$ is denoted by $s_{t}\in\mathbb{R}^{m}$ for some fixed dimension $m$ (typically $m\leq D$). We model one macro-step as
$s_{t+1}\;=\;\Psi_{t}\!\left(s_{t},\;\lambda_{t}(p)\right),\qquad t=0,1,\ldots,T_{\mathrm{macro}}-1,$ (5)
where:
  • $\Psi_{t}:\mathbb{R}^{m}\times\mathbb{R}^{r_{t}}\to\mathbb{R}^{m}$ is the ideal update map at macro-step $t$.
  • $\lambda_{t}(p)\in\mathbb{R}^{r_{t}}$ denotes the ideal parameter fragment for step $t$ encoded in the prompt.
In the construction, $\lambda_{t}(p)$ is obtained by prompt reads and simple linear post-processing, and $\Psi_{t}$ is implemented (approximately) by token-wise FFNs.
Let $\widehat{s}_{t}$ denote the corresponding state produced by the actual executor. We abstract a realized macro-step as
$\widehat{s}_{t+1}\;=\;\widehat{\Psi}_{t}\!\left(\widehat{s}_{t},\;\widehat{\lambda}_{t}(p,x)\right),\qquad t=0,1,\ldots,T_{\mathrm{macro}}-1,$ (6)
where $\widehat{\lambda}_{t}(p,x)$ is the routed-and-assembled parameter fragment (possibly depending on $x$ through the query generation), and $\widehat{\Psi}_{t}$ is the realized token-wise computation at step $t$.
E.2 Per-step error decomposition
We isolate two local error sources at step $t$.
(i) Routing error.
We assume that the assembled fragment $\widehat{\lambda}_{t}(p,x)$ is close to the ideal fragment $\lambda_{t}(p)$:
$\|\widehat{\lambda}_{t}(p,x)-\lambda_{t}(p)\|_{\infty}\;\leq\;\delta^{\mathrm{route}}_{t}\qquad\text{for all }(p,x)\in P\times X.$ (7)
In our proofs, (7) is obtained by applying the routing bounds in Appendix C to each prompt read and union-bounding over at most $R$ reads per macro-step.
(ii) Arithmetic (computation) error.
We assume that, given an ideal fragment, the realized computation approximates the ideal update uniformly:
$\sup_{s\in\mathcal{S}}\ \bigl\|\widehat{\Psi}_{t}(s,\lambda_{t}(p))-\Psi_{t}(s,\lambda_{t}(p))\bigr\|_{\infty}\;\leq\;\delta^{\mathrm{arith}}_{t},$ (8)
where $\mathcal{S}\subset\mathbb{R}^{m}$ is a bounded state set containing all intermediate states (typically $\mathcal{S}=[-B,B]^{m}$). In our proofs, (8) follows from Appendix D by approximating $\Psi_{t}(\cdot,\lambda_{t}(p))$ with a two-layer ReLU FFN to accuracy $\delta^{\mathrm{arith}}_{t}$ uniformly over $\mathcal{S}$.
(iii) Sensitivity to routed parameters.
We also record a mild Lipschitz-type sensitivity condition of the ideal update in its second argument: there exists $L_{t}\geq 0$ such that for all $s\in\mathcal{S}$ and all admissible $\lambda,\lambda^{\prime}$,
$\|\Psi_{t}(s,\lambda)-\Psi_{t}(s,\lambda^{\prime})\|_{\infty}\;\leq\;L_{t}\,\|\lambda-\lambda^{\prime}\|_{\infty}.$ (9)
In the construction, $\Psi_{t}$ is built from affine maps and bounded nonlinearities, so such an $L_{t}$ exists.
E.3 A one-step bound
Lemma E.1 (One-step error propagation).
Assume (7), (8), and (9). In addition, assume that the ideal update $\Psi_{t}$ is $K_{t}$-Lipschitz in the work-state argument on the relevant domain:
$\|\Psi_{t}(s,\lambda)-\Psi_{t}(s^{\prime},\lambda)\|_{\infty}\leq K_{t}\|s-s^{\prime}\|_{\infty}\qquad\text{for all }s,s^{\prime}\text{ and admissible }\lambda.$ (10)
Then for any $t$ and any $(p,x)\in P\times X$,
$\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}\;\leq\;K_{t}\,\|\widehat{s}_{t}-s_{t}\|_{\infty}\;+\;\delta^{\mathrm{arith}}_{t}\;+\;L_{t}\,\delta^{\mathrm{route}}_{t}.$
In particular, if $K_{t}\leq 1$ (as enforced in our executor design), this reduces to the additive bound $\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}\leq\|\widehat{s}_{t}-s_{t}\|_{\infty}+\delta^{\mathrm{arith}}_{t}+L_{t}\delta^{\mathrm{route}}_{t}$.
Proof.
Add and subtract $\Psi_{t}(\widehat{s}_{t},\lambda_{t}(p))$ and $\widehat{\Psi}_{t}(\widehat{s}_{t},\lambda_{t}(p))$:
$\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}$ $=\bigl\|\widehat{\Psi}_{t}(\widehat{s}_{t},\widehat{\lambda}_{t})-\Psi_{t}(s_{t},\lambda_{t})\bigr\|_{\infty}$
$\leq\bigl\|\widehat{\Psi}_{t}(\widehat{s}_{t},\widehat{\lambda}_{t})-\widehat{\Psi}_{t}(\widehat{s}_{t},\lambda_{t})\bigr\|_{\infty}$
$+\bigl\|\widehat{\Psi}_{t}(\widehat{s}_{t},\lambda_{t})-\Psi_{t}(\widehat{s}_{t},\lambda_{t})\bigr\|_{\infty}$
$+\bigl\|\Psi_{t}(\widehat{s}_{t},\lambda_{t})-\Psi_{t}(s_{t},\lambda_{t})\bigr\|_{\infty}.$
The middle term is bounded by $\delta^{\mathrm{arith}}_{t}$ by (8). The first term is bounded by $L_{t}\|\widehat{\lambda}_{t}-\lambda_{t}\|_{\infty}\leq L_{t}\delta^{\mathrm{route}}_{t}$ using (9) (absorbing the realized-vs-ideal dependence into the same sensitivity bound). The last term is bounded by $K_{t}\|\widehat{s}_{t}-s_{t}\|_{\infty}$ by (10). Combining yields the claim. ∎
Remark (state Lipschitz constant).
The constant $K_{t}$ in (10) depends on the ideal update $\Psi_{t}$ and the domain. If one prefers to keep the bookkeeping additive, it suffices to enforce $K_{t}\leq 1$. In our executor design this can be achieved (when needed) by scaling and residual structure, since the work-state typically enters through residual additions and coordinate-wise bounded operations.
E.4 Multi-step composition and a generic budget split
Lemma E.2 (End-to-end bound by summing local budgets).
Assume the hypotheses of Lemma E.1 for all $t=0,\ldots,T_{\mathrm{macro}}-1$ and $\widehat{s}_{0}=s_{0}$. Then for any $(p,x)\in P\times X$,
$\|\widehat{s}_{T_{\mathrm{macro}}}-s_{T_{\mathrm{macro}}}\|_{\infty}\;\leq\;\sum_{t=0}^{T_{\mathrm{macro}}-1}\left(\delta^{\mathrm{arith}}_{t}+L_{t}\,\delta^{\mathrm{route}}_{t}\right).$
Proof.
Apply Lemma E.1 iteratively and telescope using $\widehat{s}_{0}=s_{0}$. ∎
E.5 Instantiating the template
A convenient uniform allocation is to choose per-step budgets
$\delta^{\mathrm{arith}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{2T_{\mathrm{macro}}},\qquad L_{t}\,\delta^{\mathrm{route}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{2T_{\mathrm{macro}}},$
so that Lemma E.2 yields $\|\widehat{s}_{T_{\mathrm{macro}}}-s_{T_{\mathrm{macro}}}\|_{\infty}\leq\varepsilon_{\mathrm{exec}}$.
From executor state error to function error.
Let the executor output be a scalar readout $F_{\theta^{\ast}}(p,x)$ from a designated token as in Appendix B. Since the readout is affine in that token, there exists a constant $C_{\mathrm{out}}$ (e.g., $\|v_{\mathrm{out}}\|_{1}$) such that
$|F_{\theta^{\ast}}(p,x)-\widetilde{F}_{\theta^{\ast}}(p,x)|\;\leq\;C_{\mathrm{out}}\ \|\widehat{s}_{T_{\mathrm{macro}}}-s_{T_{\mathrm{macro}}}\|_{\infty},$
where $\widetilde{F}_{\theta^{\ast}}$ denotes the idealized output under perfect routing/arithmetic at the macro level. Thus, allocating budgets to make the final state error at most $\varepsilon_{\mathrm{exec}}/C_{\mathrm{out}}$ yields an end-to-end executor error at most $\varepsilon_{\mathrm{exec}}$ uniformly over $x\in X$.
Connecting to the main theorem.
In the overall proof, we first choose an MLP $N_{g}$ so that $\|N_{g}-g\|_{\infty}\leq\varepsilon_{\mathrm{approx}}$. Then we design a prompt $p_{g}$ and fixed executor $\theta^{\ast}$ so that the executor emulates $N_{g}$ with uniform error at most $\varepsilon_{\mathrm{exec}}$. By the triangle inequality,
$\sup_{x\in X}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}+\varepsilon_{\mathrm{approx}}\;\leq\;\varepsilon.$
Appendix F Prompt Encoding as Key–Value Memory
This appendix makes the ”prompt as memory” viewpoint explicit and records a concrete encoding scheme. For a prompt matrix $P=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$, define
$K_{P}:=PW_{K},\qquad V_{P}:=PW_{V},$
where $W_{K},W_{V}\in\mathbb{R}^{D\times D}$ are the same matrices used inside the attention module $\mathcal{A}_{\tau}$. Given a query $q\in\mathbb{R}^{D}$, the prompt read operation is
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}(qK_{P}^{\top})\,V_{P}\in\mathbb{R}^{D},$
where $S_{\tau}$ is the row-wise softmax with temperature $\tau>0$.
F.1 Memory layout: parameter slots and work-memory slots
In our constructions, we partition the $L$ prompt slots into parameter storage and work memory:
$L]\;=\;\mathcal{I}_{\mathrm{par}}\;\dot{\cup}\;\mathcal{I}_{\mathrm{wrk}}.$ (11)
Slots in $\mathcal{I}_{\mathrm{par}}$ store task-specific records (parameters/control signals) that the executor retrieves via $\mathcal{R}_{\tau}(\cdot;P)$. Slots in $\mathcal{I}_{\mathrm{wrk}}$ are optional scratch memory used by some variants (e.g., to cache intermediate values). In the simplest instantiation one may take $\mathcal{I}_{\mathrm{wrk}}=\emptyset$.
For each slot $j\in[L]$, we denote its projected key and value by
$k_{j}:=(K_{P})_{j,:}\in\mathbb{R}^{D},\qquad v_{j}:=(V_{P})_{j,:}\in\mathbb{R}^{D}.$
Then $\mathcal{R}_{\tau}(q;P)$ is a convex combination of the value vectors $\{v_{j}\}_{j=1}^{L}$ with weights determined by the key similarities $\{\langle q,k_{j}\rangle\}_{j=1}^{L}$.
F.2 Address labels and a fixed key codebook
Let $\mathcal{L}$ be a finite set of address labels indexing the records that may be retrieved by the executor (e.g., ”row $i$ of a matrix used at macro-step tt”, ”bias term”, etc.). Fix a key dimension $d_{k}\leq D$ and a codebook of key vectors $\{\kappa(\ell)\in\mathbb{R}^{d_{k}}:\ell\in\mathcal{L}\}$ such that there exists a separation margin $\gamma\in(0,1]$ with
[…]
For finite $\mathcal{L}$, such a codebook exists whenever $d_{k}$ is sufficiently large (e.g., using nearly-orthogonal vectors). The executor will generate a query $q(\ell)$ whose key block aligns with $\kappa(\ell)$, so that the score margin required by the routing lemma (Appendix C) holds.
Concrete choice (basis keys).
A particularly simple choice is to take $\mathcal{L}\subseteq[d_{k}]$ and set $\kappa(\ell)=e_{\ell}$, in which case (12) holds with $\gamma=1$.
F.3 Encoding labeled records into parameter slots
A record consists of a label $\ell\in\mathcal{L}$ and a payload vector $u_{\ell}\in\mathbb{R}^{D}$ (packed parameters). We implement a dictionary
$\mathcal{M}(p)\;:=\;\{(\ell,u_{\ell}):\ell\in\mathcal{L}_{\mathrm{used}}\subseteq\mathcal{L}\}$
by assigning one parameter slot $j(\ell)\in\mathcal{I}_{\mathrm{par}}$ per label $\ell$ and enforcing
$k_{j(\ell)}\approx k(\ell),\qquad v_{j(\ell)}\approx u_{\ell}.$ (13)
The approximation can be exact in the simplified executor by choosing $W_{K},W_{V}$ with a block structure that separates key-coordinates from value-coordinates (so the prompt token can independently control $k_{j(\ell)}$ and $v_{j(\ell)}$), or more generally by choosing prompt tokens $p_{j(\ell)}$ to satisfy the linear constraints induced by $W_{K},W_{V}$ on the relevant subspace.
F.4 Packing structured parameters into payload vectors
Most payloads are structured (scalars, vectors, matrix rows/blocks). We use fixed linear packing/unpacking maps
$\mathrm{pack}:\mathbb{R}^{r}\to\mathbb{R}^{D},\qquad\mathrm{unpack}:\mathbb{R}^{D}\to\mathbb{R}^{r},$
implemented inside the executor (e.g., via affine maps realizable by token-wise FFNs). For example, a scalar $\alpha\in\mathbb{R}$ may be packed as $\mathrm{pack}(\alpha)=\alpha e_{1}$, and a matrix $U\in\mathbb{R}^{a\times b}$ may be stored row-wise by packing each $U_{i,:}$ into a separate payload slot.
F.5 Retrieval guarantee (record level)
The following corollary is a direct specialization of the routing bounds in Appendix C.
Corollary F.1 (Record retrieval from parameter slots).
Fix $\tau>0$. Consider a prompt dictionary stored in slots $\{j(\ell)\in\mathcal{I}_{\mathrm{par}}\}$ with keys $\{k_{j(\ell)}\}$. Suppose that for some label $\ell^{\ast}$ and query $q=q(\ell^{\ast})$ we have the score margin
$\langle q,k_{j(\ell^{\ast})}\rangle\;\geq\;\langle q,k_{j(\ell)}\rangle+\Delta\qquad\text{for all }\ell\neq\ell^{\ast}\text{ stored in }\mathcal{M}(p).$
If all stored values satisfy $\|v_{j(\ell)}\|_{\infty}\leq B$, then
$\bigl\|\mathcal{R}_{\tau}(q;P)-v_{j(\ell^{\ast})}\bigr\|_{\infty}\;\leq\;2B\,(L-1)e^{-\Delta/\tau}.$
Proof.
Apply Lemma C.2 to the $L$ prompt slots, with $j^{\ast}=j(\ell^{\ast})$. ∎
Remark (fixed versus prompt-dependent objects).
The prompt $P$ (hence the stored keys/values) is task-dependent. All weights in the executor, including $(W_{Q},W_{K},W_{V})$, the query-generation mechanism, and the fixed packing/unpacking maps, are shared and belong to $\theta^{\ast}$.
Appendix G Executor Program: Macro-step Layout and Invariants
This appendix describes the executor as a finite “program” executed over a sequence of macro-steps. Our goals are: (i) to make explicit which subroutines are performed at each step (routing reads, arithmetic updates, optional write-backs), and (ii) to state a small set of invariants maintained throughout execution. These invariants act as the bookkeeping interface between the macro-step description and the error analysis: they specify what must hold at each step so that routing and token-wise computation can be composed with a controlled end-to-end error.
Admissible prompts (uniformity domain).
Throughout Appendix G, we restrict attention to the admissible prompt set $\mathcal{P}_{\mathrm{adm}}\subseteq\mathbb{R}^{L\times D}$ produced by the encoding scheme of Section 4.2 and Appendix F. In particular, prompts in $\mathcal{P}_{\mathrm{adm}}$ store records in fixed bounded ranges determined by the construction. All invariants and bounds in this appendix are uniform over $\mathcal{P}_{\mathrm{adm}}$ and $x\in\mathcal{X}$. See Remark 3.5 for the relation between the ambient prompt space $\mathcal{P}$ and the admissible subset $\mathcal{P}_{\mathrm{adm}}$ used for uniform bounds.
G.1 Registers and designated tokens
We consider a fixed set of designated token indices:
$\mathsf{PROMPT}:=\{1,\ldots,L\},\qquad i_{x}:=L+1,\qquad i_{\mathrm{wrk}}:=L+2,\qquad i_{\mathrm{out}}:=L+3,$
so the total sequence length is $n=L+3$. The prompt occupies $\mathsf{PROMPT}$, the input token is at $i_{x}$, the work token is at $i_{\mathrm{wrk}}$, and the output token is at $i_{\mathrm{out}}$.
We view the work token state $z^{(t)}_{i_{\mathrm{wrk}}}\in\mathbb{R}^{D}$ as a register file. Fix linear projections
$\Pi_{\mathrm{st}}:\mathbb{R}^{D}\to\mathbb{R}^{m},\qquad\Pi_{\mathrm{qry}}:\mathbb{R}^{D}\to\mathbb{R}^{D},\qquad\Pi_{\mathrm{tmp}}:\mathbb{R}^{D}\to\mathbb{R}^{m_{\mathrm{tmp}}},$
implemented by fixed affine maps in the executor (Appendix D, Lemma D.4). We interpret
$s_{t}:=\Pi_{\mathrm{st}}\!\bigl(z^{(t)}_{i_{\mathrm{wrk}}}\bigr)\in\mathbb{R}^{m}$
as the macro-state register at macro-step $t$, and
$q_{t}:=\Pi_{\mathrm{qry}}\!\bigl(z^{(t)}_{i_{\mathrm{wrk}}}\bigr)\in\mathbb{R}^{D}$
as the query register used for prompt reads. Temporary registers $\Pi_{\mathrm{tmp}}$ are used for assembling payload fragments.
G.2 Macro-step layout
We organize the depth-$T$ Transformer into $T_{\mathrm{macro}}$ macro-steps. Each macro-step consists of the following fixed pattern of subroutines:
  1. 1.
    (Query formation) From the current work token, a fixed affine (or simple FFN) subroutine writes a query vector $q_{t,r}$ into the query register for read $r\in\{1,\ldots,R_{t}\}$.
  2. 2.
    (Routing read) A single attention call reads from the prompt slots into the work token via a residual add:
    $z\leftarrow z+\mathrm{Inject}\!\bigl(\mathcal{R}_{\tau}(q_{t,r};P)\bigr),$
    where $\mathrm{Inject}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a fixed linear map selecting the coordinates/registers that receive the read vector (implemented by $W_{V}$ and subsequent fixed linear maps).
  3. 3.
    (Assembly) After $R_{t}$ reads, a fixed linear unpacking produces the assembled fragment $\widehat{\lambda}_{t}\in\mathbb{R}^{r_{t}}$ inside registers of the work token.
  4. 4.
    (Arithmetic update) A token-wise FFN applies an update to the work token so that the macro-state register $s_{t}$ is updated to $s_{t+1}$.
  5. 5.
    (Optional write-back) In variants with $\mathcal{I}_{\mathrm{wrk}}\neq\emptyset$, one additional attention call writes a bounded summary into a selected work-memory slot. (This step can be omitted in the simplest construction.)
We emphasize that the only task-dependent object is the prompt $P$ (hence $K_{P},V_{P}$). All subroutines above are realized by fixed weights in $\theta^{\ast}$.
Abstract update form.
Let $\lambda_{t}(p)$ denote the ideal parameter fragment for macro-step $t$ (as stored in $\mathcal{I}_{\mathrm{par}}$), and let $\Psi_{t}$ denote the ideal macro-update. The intended macro-dynamics is
$s_{t+1}\;=\;\Psi_{t}\!\bigl(s_{t},\lambda_{t}(p)\bigr),\qquad t=0,\ldots,T_{\mathrm{macro}}-1,$
as in Appendix E. The realized executor produces $\widehat{s}_{t}$ and $\widehat{\lambda}_{t}$ with bounded errors controlled by Appendices C and D.
G.3 Execution invariants
We now state the invariants maintained for all macro-steps $t\in\{0,\ldots,T_{\mathrm{macro}}\}$. All statements below are uniform over $P\in\mathcal{P}_{\mathrm{adm}}$ and $x\in\mathcal{X}$.
Inv-1 (Uniform boundedness).
There exists a bound $B>0$ depending only on the fixed executor architecture and construction constants (in particular on $\theta^{\ast}$, $L,D,T_{\mathrm{macro}}$, and the encoding ranges defining $\mathcal{P}_{\mathrm{adm}}$), such that all designated token vectors remain in the box $\mathcal{B}_{B}=[-B,B]^{D}$:
$\|z^{(t)}_{i}\|_{\infty}\leq B\quad\text{for all }i\in\mathsf{PROMPT}\cup\{i_{x},i_{\mathrm{wrk}},i_{\mathrm{out}}\}\text{ and all macro-times }t.$
In particular, prompt tokens satisfy $\|p_{j}\|_{\infty}\leq B$ for all $j\in[L]$, and since the value map $W_{V}$ is fixed in $\theta^{\ast}$, the prompt values $(V_{P})_{j,:}$ are also uniformly bounded (up to a constant factor absorbed into $B$), as required by the routing bound.
Inv-2 (Prompt slot partition).
The prompt slots satisfy the partition
$L]=\mathcal{I}_{\mathrm{par}}\;\dot{\cup}\;\mathcal{I}_{\mathrm{wrk}}$
as in Section 4.2, and all records needed by the program are stored in $\mathcal{I}_{\mathrm{par}}$ using the key–value encoding of Appendix F.
Inv-3 (Addressability / margin).
For each routing read $(t,r)$ there exists a target label $\ell_{t,r}$ (hence a target slot $j(\ell_{t,r})\in\mathcal{I}_{\mathrm{par}}$) such that the query $q_{t,r}$ satisfies a score margin $\Delta_{t,r}>0$:
$\langle q_{t,r},k_{j(\ell_{t,r})}\rangle\;\geq\;\langle q_{t,r},k_{j}\rangle+\Delta_{t,r}\quad\text{for all }j\in[L]\setminus\{j(\ell_{t,r})\}.$
Inv-4 (Register integrity).
Each subroutine only writes to intended registers: query formation only changes $\Pi_{\mathrm{qry}}$, routing injection only affects designated payload registers, and the arithmetic FFN updates $\Pi_{\mathrm{st}}$ according to the intended $\Psi_{t}$ while preserving boundedness (consistent with Inv-1).
G.4 Maintaining Inv-1: a generic sufficient condition
The next lemma records a clean sufficient condition under which the uniform boundedness invariant (Inv-1) holds. Its role is to align the boundedness bookkeeping with the actual executor dynamics (which includes residual additions).
Lemma G.1 (Sufficient condition for Inv-1).
Assume that the construction fixes $B_{0}>0$ such that:
  1. (i)
    The encoding maps used to form the initial tokens $(P,u(x),w,o)$ produce vectors in $\mathcal{B}_{B_{0}}$ uniformly over $x\in\mathcal{X}$ and all $P\in\mathcal{P}_{\mathrm{adm}}$.
  2. (ii)
    Each macro-step subroutine (query formation, routing injection, linear assembly, and arithmetic FFN) is implemented by fixed weights chosen so that $\mathcal{B}_{B_{0}}$ is an invariant set for the executor dynamics, i.e., each subroutine maps inputs in $\mathcal{B}_{B_{0}}$ to outputs in $\mathcal{B}_{B_{0}}$. (For instance, this can be ensured by range-restricted encodings together with appropriate scaling of linear maps so that residual additions remain within $\mathcal{B}_{B_{0}}$ on that set; explicit clipping/saturation is one sufficient but not necessary mechanism.)
Then Inv-1 holds with $B=B_{0}$.
Proof.
By induction over macro-steps. The base case follows from (i). For the inductive step, assume all designated tokens lie in $\mathcal{B}_{B_{0}}$ at macro-time $t$. Applying the subroutines of macro-step $t$ preserves membership in $\mathcal{B}_{B_{0}}$ by (ii), hence the designated tokens remain bounded at macro-time $t+1$. ∎
G.5 Maintaining invariants and deriving per-step bounds
The next lemma shows that Inv-3 implies an exponentially small routing error for each read, and that Inv-1 guarantees that the arithmetic approximation can be taken uniformly on a fixed compact domain.
Lemma G.2 (Per-read routing and per-step arithmetic prerequisites).
Assume Inv-1–Inv-3 hold at macro-step $t$. Let $q_{t,r}$ be a query with margin $\Delta_{t,r}$ for its target slot. Then the routed read obeys
$\bigl\|\mathcal{R}_{\tau}(q_{t,r};P)-(V_{P})_{j(\ell_{t,r}),:}\bigr\|_{\infty}\;\leq\;2B\,(L-1)e^{-\Delta_{t,r}/\tau}.$
Moreover, by Inv-1, all token-wise FFN approximations invoked in the arithmetic step can be taken uniformly over $\mathcal{B}_{B}=[-B,B]^{D}$.
Proof.
The routing bound is exactly Lemma C.2 (Appendix C). The arithmetic prerequisite follows since Inv-1 ensures that all FFN inputs remain in $\mathcal{B}_{B}$, the compact domain on which Appendix D provides uniform approximation. ∎
G.6 A generic macro-step error statement
We now state the executor-level conclusion in a form that plugs into the composition template.
Proposition G.3 (Macro-step realization with budgeted errors).
Fix macro-step $t$ and suppose Inv-1–Inv-4 hold. Assume that:
  • Each routing read $(t,r)$ satisfies a margin $\Delta_{t,r}$ and uses temperature $\tau$, so that the total assembled parameter fragment $\widehat{\lambda}_{t}$ obeys
    $\|\widehat{\lambda}_{t}-\lambda_{t}(p)\|_{\infty}\leq\delta^{\mathrm{route}}_{t}$
    for some prescribed $\delta^{\mathrm{route}}_{t}$ (obtained by summing the per-read bounds and the linear assembly operator norm).
  • The token-wise FFN used in the arithmetic step approximates the ideal macro-update $(s,\lambda)\mapsto\Psi_{t}(s,\lambda)$ to accuracy $\delta^{\mathrm{arith}}_{t}$ uniformly over the bounded state set $\mathcal{S}\subseteq\mathbb{R}^{m}$ and admissible $\lambda$.
Then the work-state register satisfies the one-step bound
$\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}\;\leq\;\|\widehat{s}_{t}-s_{t}\|_{\infty}\;+\;\delta^{\mathrm{arith}}_{t}\;+\;L_{t}\,\delta^{\mathrm{route}}_{t},$
where $L_{t}$ is a Lipschitz constant of $\Psi_{t}$ in its parameter argument (as in Appendix E).
Proof.
This is a direct instantiation of Lemma E.1 in Appendix E. The only additional ingredient is that the realized fragment $\widehat{\lambda}_{t}$ is produced by finitely many reads followed by fixed linear assembly; hence its deviation from $\lambda_{t}(p)$ is bounded by the sum of the individual routing errors multiplied by the assembly operator norm. ∎
G.7 End-to-end view
Combining Proposition G.3 over $t=0,\ldots,T_{\mathrm{macro}}-1$ and using the budget allocation in Appendix E yields a uniform executor emulation bound
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p,x)-\widetilde{F}_{\theta^{\ast}}(p,x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}},$
where $\widetilde{F}_{\theta^{\ast}}$ denotes the idealized program output under perfect routing and perfect arithmetic. This is the executor-side half of the triangle-inequality split used in the main proof.
Remark (what must be specified in the main construction).
To fully instantiate this appendix for the final proof, the main construction must specify: (i) the label set $\{\ell_{t,r}\}$ and memory layout in $\mathcal{I}_{\mathrm{par}}$, (ii) how queries $q_{t,r}$ are generated and why they satisfy the margin condition (Inv-3), and (iii) why the uniform boundedness (Inv-1) holds with a constant $B$ uniform over all admissible prompts in $\mathcal{P}_{\mathrm{adm}}$, under the chosen encodings, injections, and arithmetic updates. These are precisely the points addressed by the routing, encoding, arithmetic, and composition appendices.
Appendix H Full Proof of the Main Theorem
In this appendix we give a complete proof of the main theorem stated in the main text. This section turns the proof sketch into a fully detailed argument by explicitly carrying out the approximation split, the prompt-encoding step, and the executor emulation bound, and then combining them into the final uniform error guarantee.
Theorem H.1 (Main theorem (prompt switching with a fixed executor)).
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be compact as in Assumption 3.1. There exists a fixed set of executor parameters $\theta^{\ast}$ (a simplified Transformer as in Appendix B) such that for every $g\in C(\mathcal{X})$ and every $\varepsilon>0$ there exists a soft prompt $p_{g}\in\mathcal{P}$ with
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;<\;\varepsilon.$
Equivalently, for the prompt-induced family $H_{\theta^{\ast}}=\{F_{\theta^{\ast}}(p,\cdot):p\in\mathcal{P}\}$ (Definition 3.4), we have $\overline{H_{\theta^{\ast}}}^{\|\cdot\|_{\infty}}=C(\mathcal{X})$.
Proof.
Fix an arbitrary target $g\in C(\mathcal{X})$ and $\varepsilon>0$. Set
$\varepsilon_{\mathrm{approx}}:=\varepsilon/2,\qquad\varepsilon_{\mathrm{exec}}:=\varepsilon/2,\qquad\text{so that }\ \varepsilon_{\mathrm{approx}}+\varepsilon_{\mathrm{exec}}=\varepsilon.$
Step 1: Classical approximation by a small ReLU MLP.
By the classical universal approximation theorem for ReLU networks on compact domains, there exists a (finite) ReLU MLP $N_{g}:\mathcal{X}\to\mathbb{R}$ such that
$\sup_{x\in\mathcal{X}}|N_{g}(x)-g(x)|\;\leq\;\varepsilon_{\mathrm{approx}}.$ (14)
Without loss of generality, choose $N_{g}$ to be a depth-$S=3$ (one-hidden-layer) ReLU MLP in the sense of Definition 7.1. In particular, write
$N_{g}(x)=a^{\top}\sigma(Wx+b)+c\quad\text{and identify}\quad W^{(1)}=W,\ b^{(1)}=b,\ W^{(2)}=a^{\top},\ b^{(2)}=c.$
Collect the parameters in the same format as Definition 7.1:
$\pi_{g}:=\big(W^{(1)},b^{(1)},W^{(2)},b^{(2)}\big),$
and define the (finite) bound
$\Lambda:=\max\Big\{\|W^{(1)}\|_{\max},\|b^{(1)}\|_{\max},\|W^{(2)}\|_{\max},|b^{(2)}|\Big\}<\infty.$
Thus $N_{g}$ belongs to the bounded class in Definition 7.1 for this $(g,\varepsilon)$-dependent $\Lambda$, and (14) holds.
Step 2: Fix a single executor $\theta^{\ast}$ (independent of $g$).
We now fix $\theta^{\ast}$ once and for all, using the simplified Transformer architecture in Appendix B and the macro-step ”program” layout in Appendix G. Concretely, $\theta^{\ast}$ contains: (i) fixed attention and FFN weights realizing the routing and arithmetic subroutines, (ii) fixed linear packing/unpacking maps used to assemble parameter fragments, (iii) a fixed output readout map producing $F_{\theta^{\ast}}(p,x)$ from a designated token. No part of $\theta^{\ast}$ depends on $g$.
The executor program is designed so that, under ideal reads and ideal token-wise computations, it performs the forward pass of any target MLP whose parameters are presented to it through prompt memory. This is formalized by the macro-step semantics and invariants in Appendix G.
Step 3: Encode the MLP parameters into a prompt.
Using the key–value prompt encoding of Appendix F, we store the finite parameter list $\pi_{g}$ into the prompt as a dictionary of labeled records in the parameter slot set $\mathcal{I}_{\mathrm{par}}$, with the slot partition
$L]=\mathcal{I}_{\mathrm{par}}\;\dot{\cup}\;\mathcal{I}_{\mathrm{wrk}}.$ (15)
Each record label specifies which fragment is needed at which macro-step (e.g., ”layer $\ell$ weight row ii”, ”layer $\ell$ bias block”), and each payload is produced by a fixed packing map. This yields a prompt matrix $p_{g}\in\mathbb{R}^{L\times D}$ (equivalently, $p_{g}\in\mathcal{P}$).
Step 4: Choose routing and arithmetic accuracies and verify emulation.
Let $T_{\mathrm{macro}}$ be the number of macro-steps in the executor program, and let $R$ be an upper bound on the number of prompt reads per macro-step (both are fixed by $\theta^{\ast}$ and do not depend on $g$). We allocate a uniform local error budget by choosing per-macro-step tolerances
$\delta^{\mathrm{arith}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{2T_{\mathrm{macro}}},\qquad L_{t}\,\delta^{\mathrm{route}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{2T_{\mathrm{macro}}},\qquad t=0,\ldots,T_{\mathrm{macro}}-1,$ (16)
where $L_{t}$ is a (fixed) Lipschitz sensitivity constant for the ideal macro-update at step $t$ as in Appendix E. (If one prefers, one may upper bound all $L_{t}$ by a single constant $L_{\max}$ and allocate using $L_{\max}$.)
Routing. For each read in each macro-step, the executor forms a query $q$ targeting the corresponding stored label. By the margin invariant in Appendix G and the routing bound in Appendix C, we can choose a temperature $\tau$ (and, if needed, scaling inside the fixed query generation) so that every read has impurity at most $\rho$, hence each read error is at most $2B\rho$ in $\|\cdot\|_{\infty}$ for a uniform value bound $B$. Since each macro-step uses at most $R$ reads and the subsequent assembly is a fixed linear map, we can choose $\rho$ (equivalently $\tau$) small enough so that the assembled parameter error satisfies
$\|\widehat{\lambda}_{t}(p_{g},x)-\lambda_{t}(p_{g})\|_{\infty}\;\leq\;\delta^{\mathrm{route}}_{t}\qquad\text{uniformly over }x\in\mathcal{X}.$ (17)
Arithmetic. On the bounded state set guaranteed by the boundedness invariant in Appendix G, the token-wise two-layer ReLU FFN can approximate the ideal macro-update map uniformly to accuracy $\delta^{\mathrm{arith}}_{t}$ by Appendix D. Thus the realized arithmetic subroutine satisfies
$\sup_{x\in\mathcal{X}}\bigl\|\widehat{\Psi}_{t}(\widehat{s}_{t},\lambda_{t}(p_{g}))-\Psi_{t}(\widehat{s}_{t},\lambda_{t}(p_{g}))\bigr\|_{\infty}\;\leq\;\delta^{\mathrm{arith}}_{t}.$ (18)
Combining (17) and (18) across all macro-steps using the composition template in Appendix E (Lemma E.2) yields
$\sup_{x\in\mathcal{X}}\ \|\widehat{s}_{T_{\mathrm{macro}}}(x)-s_{T_{\mathrm{macro}}}(x)\|_{\infty}\;\leq\;\sum_{t=0}^{T_{\mathrm{macro}}-1}\left(\delta^{\mathrm{arith}}_{t}+L_{t}\,\delta^{\mathrm{route}}_{t}\right)\;\leq\;\varepsilon_{\mathrm{exec}}.$ (19)
Finally, since the scalar readout $F_{\theta^{\ast}}(p_{g},x)$ is an affine function of a designated token/register (Appendix B), the state error bound (19) implies an output error bound
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-N_{g}(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}.$ (20)
(Any readout Lipschitz constant can be absorbed by tightening the local budgets in (16).)
Step 5: Conclude by the triangle inequality.
By (14) and (20),
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;\leq\;\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-N_{g}(x)\bigr|+\sup_{x\in\mathcal{X}}\bigl|N_{g}(x)-g(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}+\varepsilon_{\mathrm{approx}}\;=\;\varepsilon.$
Since $g\in C(\mathcal{X})$ and $\varepsilon>0$ were arbitrary, the theorem follows. ∎